### v1

# Nonlinear semiclassical spectroscopy of ultrafast molecular polariton dynamics

Abstract
We introduce a theoretical framework that allows for the systematic and efficient description of the ultrafast nonlinear response of molecular polaritons, i.e., hybrid light-matter states, in the collective regime of large numbers of molecules $\mathcal{N}$ coupled to the cavity photon mode. Our approach is based on a semiclassical, mean-field evolution of the molecular Hamiltonian and the cavity field, which is complemented by a perturbative expansion of both light and matter counterparts in the input fields entering the cavity. In addition, expansion in terms of the pulse phases enables us to disentangle different excitation pathways in Liouville space, thereby distinguishing contributions to the nonlinear response. The formalism extends traditional free-space nonlinear spectroscopy by incorporating the feedback of matter onto the light field via the induced polarization. We demonstrate the utility of the framework by applying it to the calculation of pump-probe polariton spectra and show how, by storing the pulses, the cavity facilitates additional excitation pathways which are not possible in free space. Our method, which does not scale with $\mathcal{N}$, is broadly applicable and can be extended to a wide range of current experiments investigating the dynamical nonlinear response of hybrid light-matter states.
pacs:
42.50.Pq, 42.65.-k, 42.65.Sf
Figure 1: (a) Schematics. An optical cavity containing an ensemble of $\mathcal{N}$ molecules is driven by input pulses with amplitudes $\eta_{p}$ (pump) and $\eta_{p^{\prime}}$ (probe). (b) Key distinction between conventional, free-space nonlinear spectroscopy and nonlinear cavity spectroscopy, concerning the evolution of the molecular density matrix. For cavity spectroscopy, the feedback of the molecular polarization onto the cavity field leads to a coupling of the $n$-th order molecular density matrix $\vec{\rho}^{(n)}$ to all lower orders in the cavity field, as opposed to the free-space case, where the interaction is always linear in the laser field that drives the molecules. (c) Pump and probe pulses entering the cavity with time delay $\tau_{\Delta}$ get stored and prolonged in the cavity on a time scale given by the cavity lifetime $\kappa^{-1}$, therefore allowing the cavity field created by the probe ($\alpha^{(0)(1)}$) to act before the pump ($\alpha^{(1)(0)}$). Additional phases $\Phi_{p,p^{\prime}}$ imprinted on the pulses can be utilized to separate excitation pathways in Liouville space.
Introduction.—Molecular polaritons, quasiparticles resulting from the strong coupling between confined photonic modes and vibrational or electronic excitations of molecules, e.g., inside Fabry-Pérot cavities, have gained significant attention over recent years, largely due to their potential applications in areas such as polariton chemistry Hutchison et al. (2012); Thomas et al. (2019), energy transport Coles et al. (2014); Zhong et al. (2017), condensation and lasing Kéna-Cohen and Forrest (2010); Plumhof et al. (2014), and the manipulation of optical nonlinearities Barachati et al. (2018); Xiang et al. (2019); Wang et al. (2021); Cheng et al. (2022). Recent advances in electronic Takemura et al. (2015a, b); DelPo et al. (2020); Fassioli et al. (2021); Mewes et al. (2020); Son et al. (2022); Wu et al. (2022); Russo et al. (2024) and vibrational Xiang et al. (2018, 2019); Xiong (2023); Sufrin et al. (2024); Dunkelberger et al. (2022); Duan et al. (2021); Simpkins et al. (2023) nonlinear polariton spectroscopy have provided new insight into the nonlinear response of molecular polaritons by revealing the dynamics of the complex interactions and energy relaxation pathways. However, significant challenges and open questions remain in understanding the nonlinear polariton response, particularly regarding the short-time dynamical behavior, i.e., before the cavity photons, and thus the polariton states, have decayed ($\sim$ fs in experiments). The development of an efficient formalism that can address these questions is crucial for advancing the understanding of polariton dynamics across different time scales as observed in pump-probe experiments.
While the theoretical formalism of nonlinear spectroscopy involving the interaction between laser light and matter in free space is well-established Mukamel (1995); Yuen-Zhou et al. (2014); Jonas (2003); Cho (2008); Gelin et al. (2009), recent works have aimed at constructing a framework for the theory of of nonlinear electronic and vibrational polariton spectroscopy Saurabh and Mukamel (2016); F. Ribeiro et al. (2018); Zhang et al. (2023); Mondal et al. (2023); Gallego-Valencia et al. (2024); Schnappinger et al. (2024). Some studies have considered direct excitation of the molecules through the side of the cavity Zhang et al. (2023), whereas other works have taken a fully quantum mechanical approach, thereby restricting the scope of the treatment to either long times after light and matter have decoupled, or the explicit simulation of only a few molecules F. Ribeiro et al. (2018); Gallego-Valencia et al. (2024).
In this Letter, we introduce a systematic formalism to nonlinear cavity spectroscopy in the limit of large numbers of molecules $\mathcal{N}$ per photon mode. Our approach is based on a mean-field evolution of the coupled light-matter system Fowler-Wright et al. (2022), related to semiclassical Maxwell-Liouville approaches which have been developed, e.g., in semiconductor optics and nanophotonics Jahnke et al. (1996); Scully and Zubairy (1997); Lopata and Neuhauser (2009); Šindelka (2010); Sukharev and Nitzan (2011); Li et al. (2018); Chen et al. (2019). We perturbatively expand both cavity field and density matrix in the input fields driving the cavity and show that the formalism can be regarded as an extension of the conventional free-space perturbative nonlinear spectroscopy framework by Mukamel Mukamel (1995), now incorporating nonlinear contributions to the field resulting from the feedback of the molecules onto the cavity dynamics. In addition, expansion of the nonlinear response in terms of phase components allows us to separate different excitation pathways in Liouville space and identify new pathways enabled by the cavity, using the widely-adopted toolkit of double-sided Feynman diagrams.
Mean-field approach.—We consider an ensemble of $\mathcal{N}$ molecules collectively coupled to the single confined mode of an optical cavity at frequency $\omega_{c}$ [see schematics in Fig. 1(a)]. Each molecule $j$ is described by an (in principle arbitrary) Hamiltonian $\mathcal{H}_{0}^{j}$ with a corresponding dipole operator $\hat{\mu}_{j}$. The interaction between molecules and the cavity field is described by a Dicke-type Hamiltonian such that the total Hamiltonian describing the full quantum evolution of the coupled light-matter system expresses as (we set $\hbar\equiv 1$)
$\mathcal{H}=\omega_{c}\hat{a}^{\dagger}\hat{a}+\sum_{j=1}^{\mathcal{N}}\mathcal{H}_{0}^{j}+E_{0}(\hat{a}+\hat{a}^{\dagger})\sum_{j=1}^{\mathcal{N}}\hat{\mu}_{j},$ (1)
with the cavity zero-point amplitude $E_{0}=\sqrt{\omega_{c}/(2\epsilon_{0}\mathcal{V})}$, where $\mathcal{V}$ is the cavity mode volume and $\epsilon_{0}$ is the vacuum permittivity. To facilitate the understanding of the formalism, we initially assume that the cavity field is driven by only a single input field with amplitude $\eta$ at a central carrier frequency $\omega_{\ell}$ and with a Gaussian pulse shape described by the temporal envelope function $f(t)$ with a temporal pulse width $\tau_{w}$.
The limit of large $\mathcal{N}$ allows us to apply a mean-field approximation (exact for $\mathcal{N}\to\infty$), which amounts to setting $\expectationvalue{\hat{a}}=\alpha(t)$ Fowler-Wright et al. (2022). Additionally, for illustration, we take all molecules to be identical, i.e., $\hat{\mu}_{j}=\hat{\mu}$, although extension to disordered ensembles offers no conceptual difficulty. The simplified mean-field Hamiltonian then expresses as
$\mathcal{H}_{\text{MF}}=\mathcal{H}_{0}+E_{0}\left(\alpha(t)+\alpha^{*}(t)\right)\hat{\mu}.$ (2)
The dynamical evolution of the (now single-particle) molecular density matrix $\rho$ is then governed by the master equation $\partial_{t}\rho=\mathrm{i}[\rho,\mathcal{H}_{\text{MF}}]+\mathcal{D}[\rho]$. Here, the dissipator $\mathcal{D}[\rho]$ represents possible additional loss channels of the molecule (e.g., dephasing at rate $\gamma_{\phi}$, non-radiative decay,$\ldots$) that may be added, e.g., in Lindblad form. However, a complete description of the mean-field problem also involves solving the classical equation for the cavity field amplitude which undergoes damping at rate $\kappa$ (photon loss)
$\dot{\alpha}(t)$ $=-\left(\frac{\kappa}{2}+\mathrm{i}\omega_{c}\right)\alpha(t)-\mathrm{i}\mathcal{N}E_{0}P(t)-\eta f(t)\mathrm{e}^{-\mathrm{i}\omega_{\ell}t},$ (3)
containing the induced molecular polarization $P(t)=\mathrm{Tr}[\hat{\mu}\rho(t)]$ created by all $\mathcal{N}$ molecules. Altogether, this describes a self-consistent, semiclassical evolution of the coupled light-matter system, where the matter component evolves quantum mechanically, while the part describing the electromagnetic field undergoes classical evolution and is driven by both the input field as well as by the collective molecular polarization feeding back into the cavity field.
Perturbative expansion.—While the above set of mean-field equations can be readily solved numerically, there is generally no closed analytical solution due to the nonlinearity of the molecular transitions. To gain more insight into the structure of the nonlinear response, we therefore proceed in analogy to standard nonlinear spectroscopy with a perturbative expansion of both the field and the matter density matrix in terms of the input field amplitude, i.e.,
$\alpha(t)=\sum_{n=0}^{\infty}\eta^{n}\alpha^{(n)}(t),\quad\rho(t)=\sum_{n=0}^{\infty}\eta^{n}\rho^{(n)}(t).$ (4)
From this, also other perturbative quantities can be derived such as, e.g., the $n$-th order polarization $P^{(n)}(t)=\mathrm{Tr}[\hat{\mu}\rho^{(n)}(t)]$  [^1] Note that the $\rho^{(n)}$ are not true density matrices, i.e., they do not obey the properties of density matrices such as the conservation of the trace.
Furthermore, for both numerical and analytical considerations, it is advantageous to convert the master equation for the density matrix onto matrix-vector form, i.e., map the density matrix $\rho$ onto a vector $\vec{\rho}$ and the superoperator acting on the density matrix onto a matrix (also referred to as Liouville space) [see Supplementary Material (SM) sm for details on vectorization] Mukamel (1995); Horn and Johnson (1994); Am-Shallem et al. (2015). Then, one obtains a closed system for the cavity field amplitude and the vectorized density matrix up to $n$-th order
$\dot{\alpha}^{(n)}$ $=-\left(\frac{\kappa}{2}+\mathrm{i}\omega_{c}\right)\alpha^{(n)}-\mathrm{i}\mathcal{N}E_{0}P^{(n)}-\delta_{1n}f(t)\mathrm{e}^{-\mathrm{i}\omega_{\ell}t},$ (5a)
$\dot{\vec{\rho}}^{(n)}$ $=-\mathrm{i}\mathcal{L}_{0}\vec{\rho}^{(n)}-\mathrm{i}E_{0}\sum_{j=0}^{n}\left(\alpha^{(n-j)}+\alpha^{(n-j)*}\right)\mathcal{L}_{\mu}\vec{\rho}^{(j)},$ (5b)
in terms of the vectorized commutators $[\mathcal{H}_{0},\rho^{(n)}]\to\mathcal{L}_{0}\vec{\rho}^{(n)}$ and $[\hat{\mu},\rho^{(j)}]\to\mathcal{L}_{\mu}\vec{\rho}^{(j)}$ where $\mathcal{L}_{0}$, $\mathcal{L}_{\mu}$ are the free molecular and interaction Liouvillians, respectively. In particular, Eqs. (5) indicate that only the first-order cavity field is driven directly by the input fields while for higher orders, the molecular density matrix is driven by lower orders of the cavity field. Importantly, Eq. (5b) shows that the (cavity) field driving the molecular density matrix can have nonlinear contributions in the input field, as opposed to free-space spectroscopy where the interaction Liouvillian driving the evolution of the density matrix is always linear in the (laser) input field [see Fig. 1(b)] Mukamel (1995). In other words, the electric field experienced by the molecules is not the laser field interacting with the cavity from outside, but a distorted version of it due to the evolving material polarization inside the cavity. While for conventional spectroscopy with well-separated pulses, only a single multidimensional integral is obtained as solution for the $n$-th order density matrix, many additional integrals arise in the cavity scenario ($12^{n-1}-$ 1, see diagram in Fig. 2 and SM sm ). Also, a straightforward analysis of Eqs. (5) reveals that the first order ($n=1$) recovers the standard linear polariton response, characterized by a Rabi splitting proportional to $\sqrt{\mathcal{N}}$, arising from the hybridization of the cavity with the collective molecular bright mode Rocca et al. (1998); Agranovich et al. (2003); Litinskaya et al. (2004); Gonzalez-Ballestero et al. (2016); Campos-Gonzalez-Angulo and Yuen-Zhou (2022); Yuen-Zhou and Koner (2024).
Figure 2: Density matrix evolution. Graph illustrating the possible integrals contributing to the dynamical evolution of the $n$-th order density matrix $\rho^{(n)}(t)$. The orange path on the right illustrates the integral that is obtained for conventional free-space spectroscopy, corresponding to a sequence of first-order interactions linear in the input. The feedback of the nonlinear polarization onto the cavity field adds $12^{n-1}-$ 1 other possible pathways.
Figure 3: Differential transmission. DT spectra $\Delta T(\omega)$ for a 2LS with resonance frequency $\omega_{0}$ as a function of the delay time between the pump and probe pulse $\tau_{\Delta}$ for (a) weak ($\gamma_{\phi}=0$), and (b) strong ($\gamma_{\phi}=\kappa$) dephasing of the electronic transition. The green dashed lines show the location of the (linear response) polariton frequencies. The black curve in (b) shows a cross section through the DT at $\tau_{\Delta}=4\kappa^{-1}$. The plots in (c) and (d) show the dynamics of the population and polarization $|P(t)|^{2}$ created by the initial pump pulse arriving at $\tau_{p}=2\kappa^{-1}$ which are then probed by the second pulse after $\tau_{\Delta}$. The 2LS is initialized in the ground state, with parameters $g\sqrt{\mathcal{N}}=3\kappa$, $\tau_{w}=0.05\kappa^{-1}$, $\omega_{0}=\omega_{c}$, and central frequency $\omega_{p}=\omega_{p^{\prime}}=\omega_{0}$ for both pulses.
Differential transmission.—We now consider two pulses driving the cavity (pump $p$ and probe $p^{\prime}$), with input fields proportional to $\eta_{j}f_{j}(t-\tau_{j})\mathrm{e}^{-\mathrm{i}\omega_{j}t}$ where $\eta_{j}$, $f_{j}$, $\omega_{j}$, $\tau_{j}$ for $j\in\{p,p^{\prime}\}$ denote the amplitude, envelope, carrier frequency, and arrival time of the pump and probe pulses, respectively. We assume $\tau_{p^{\prime}}>\tau_{p}$, with a time delay between pump and probe of $\tau_{\Delta}=\tau_{p^{\prime}}-\tau_{p}$ [see sketch in Fig. 1(c)]. We proceed with a perturbative expansion in both pump ($p$) and probe ($p^{\prime}$) amplitudes (this can be generalized to arbitrary many pulses)
$\alpha(t)$ $=\sum_{n,m=0}^{\infty}\eta_{p}^{n}\eta_{p^{\prime}}^{m}\alpha^{(n)(m)}(t),$ (6a)
$\rho(t)$ $=\sum_{n,m=0}^{\infty}\eta_{p}^{n}\eta_{p^{\prime}}^{m}\rho^{(n)(m)}(t),$ (6b)
where the first and second indeces refer to the orders in the pump and probe fields, respectively. This allows us to derive a perturbative set of equations for $\alpha^{(n)(m)}$, $\vec{\rho}^{(n)(m)}$ similar to Eqs. (5), now describing pump-probe spectroscopy (see SM sm for details).
Figure 4: Separating phase contributions to the nonlinear polariton response. (a), (b) Plot of phase contributions $(0,1)$ and $(2,-1)$ to $\Delta T(\omega)$ as a function of the delay time between the pulses. The $(2,-1)$-contribution decays on a time scale of $\kappa^{-1}$, as it relies on the probe acting prior to the pump. The green dashed lines show the (linear response) polariton frequencies. We chose $g\sqrt{\mathcal{N}}=3\kappa$, $\gamma_{\phi}=0.1\kappa$, $\omega_{p}=\omega_{p^{\prime}}=\omega_{0}$, and a temporal width of $\tau_{w}=0.1\kappa^{-1}$ for both pulses. The double-sided Feynman diagrams in (c) and (d) represent processes contributing to the different phase combinations of $\alpha^{(2)(1)}$. The dotted arrows indicate the decay of the coherence.
From the solution for the total cavity field, the cavity transmission and reflection can be obtained using standard input-output theory Steck (2007). Since nonlinearities are typically weak, instead of absolute quantities, differential (nonlinear) corrections to the response are commonly considered. For the collinear geometry in question, we define the differential transmission (DT) spectrum by subtracting the (normalized) probe transmission without the pump from the probe transmission in the presence of the pump pulse
$\Delta T(\omega)=T_{p^{\prime}}^{\text{pump\, on}}(\omega)-T_{p^{\prime}}^{\text{pump\, off}}(\omega),$ (7)
where the transmission is related to the intracavity field $T_{p^{\prime}}=|(\kappa/2)\alpha(\omega)/[\eta_{p^{\prime}}f_{p^{\prime}}(\omega)]|^{2}$. The contribution $T_{p^{\prime}}^{\text{pump\, on}}(\omega)$ depends on the time delay between the pulses $\tau_{\Delta}$. By considering only the lowest (third) order correction due to pump and probe, the DT may be expressed as (see SM for derivation sm )
$\Delta T(\omega)\approx\left(\frac{\kappa}{2}\right)^{2}\eta_{p}^{2}\frac{2\mathrm{Re}\left[\alpha^{(0)(1)*}(\omega)\alpha^{(2)(1)}(\omega)\right]}{f_{p^{\prime}}(\omega)^{2}},$ (8)
where the key quantity is the third-order cavity field $\alpha^{(2)(1)}$. Beware that Eq. (8) is not the usual pump-probe signal; instead, it is a signal which contains different phase contributions (see paragraph below). The DT is plotted as a function of delay time in Figs. 3(a), (b) for two-level systems (2LSs) with transition dipole moment $\mu_{ge}$ coupled to the cavity at strength $g=\mu_{ge}E_{0}$, for various degrees of dephasing of the transition. The plots below in Figs. 3(c), (d) show the dynamics of the polarization and population created by an initial pump pulse which are then probed after a time $\tau_{\Delta}$. At early times, oscillatory behavior around the polariton frequencies is observed, corresponding to coherences between upper and lower polariton. Dephasing leads to a decay of coherence and the creation of stationary population in the molecules, even after the cavity photon has decayed. In the long-time limit, this manifests as a contracted Rabi splitting, characterized by a DT as shown by the black cross-section curve in Fig. 3(b) Xiang et al. (2018); Dunkelberger et al. (2022); F. Ribeiro et al. (2018).
Separating excitation pathways.—The nonlinear signal stemming from the $n$-th order density matrix generally contains many possible excitation pathways. For instance, even for free-space spectroscopy of a 2LS, there are eight Feynman diagrams contributing to the third-order nonlinear response Mukamel (1995). In experiments, it is therefore highly desired to isolate certain excitation pathways and thereby understand the different contributions to the nonlinear signal, which, owing to the collinear geometry in question, can be achieved by phase cycling. This method is equivalent to the more standard phase matching procedure, which, however obfuscates our formalism as it requires the description of a multimode cavity Yuen-Zhou et al. (2014); Jonas (2003); Hybl et al. (1998); Tian et al. (2003); Tan (2008); Yang et al. (2020).
To this end, we will consider the effect of additional phases applied to the pump ($\Phi_{p}$) and probe ($\Phi_{p^{\prime}}$) pulses. These phases are subsequently also imprinted onto the cavity field and the molecular density matrix, enabling us to disentangle different excitation pathways characterized by specific phase sequences Mukamel (1995); Yuen-Zhou et al. (2014). The cavity field and density matrix are now expanded as
$\alpha(t)$ $=\sum_{n,m,\vec{v}}\eta_{p}^{n}\eta_{p^{\prime}}^{m}\mathrm{e}^{-\mathrm{i}\vec{v}\cdot\vec{\Phi}}\alpha^{(n)(m)}_{\vec{v}}(t),$ (9a)
$\rho(t)$ $=\sum_{n,m,\vec{v}}\eta_{p}^{n}\eta_{p^{\prime}}^{m}\mathrm{e}^{-\mathrm{i}\vec{v}\cdot\vec{\Phi}}\rho^{(n)(m)}_{\vec{v}}(t),$ (9b)
where $\vec{\Phi}=(\Phi_{p},\Phi_{p^{\prime}})^{\top}$ denotes the vector of pump and probe phases, and $\vec{v}=(v_{p},v_{p^{\prime}})^{\top}$ with $v_{p},v_{p^{\prime}}\in\mathbb{Z}_{0}$ denotes the phase coefficients which also depend on the pump and probe orders $v_{p}(n)$, $v_{p^{\prime}}(m)$ (see SM sm ). The raw DT signal from Eq. (8) can then be separated into different phase contributions
$\Delta T_{\vec{v}}(\omega)\approx\left(\frac{\kappa}{2}\right)^{2}\eta_{p}^{2}\frac{2\mathrm{Re}\left[\alpha^{(0)(1)*}_{(0,-1)}(\omega)\alpha^{(2)(1)}_{\vec{v}}(\omega)\right]}{f_{p^{\prime}}(\omega)^{2}},$ (10)
with the total signal simply being obtained as the sum of all contributions $\Delta T=\sum_{\vec{v}}\Delta T_{\vec{v}}$. One finds that there are two possible phase combinations contributing to $\alpha^{(2)(1)}$: $(0,1)$, and $(2,-1)$ [and the complex conjugate processes $(0,-1)$, $(-2,1)$], which are shown in Figs. 4(a), (b), and are associated with pump-probe and double-quantum coherence (DQC) spectroscopies, respectively. The first contribution [$(0,1)\equiv\pm\Phi_{p}\mp\Phi_{p}+\Phi_{p^{\prime}}$] is the ‘standard’ pump-probe, where populations and zeroth-order quantum coherences are created by twofold application of the pump with opposite phase [see Feynman diagrams in Fig. 4(c)]. The coherent superposition between upper and lower polaritons then undergo Rabi oscillations which are detected by the probe. The second phase combination [($(2,-1)\equiv 2\Phi_{p}-\Phi_{p^{\prime}}$] is negligible outside of the cavity, as it implies that the 2LS is excited twice. However, inside the cavity, since the pulses get extended on a time scale of $\kappa^{-1}$ [cf. Fig. 1(c)], the probe can act before the pump, therefore enabling this phase combination to arise within $\kappa^{-1}$. In the SM sm , we derive the phase contributions up to third order for an arbitrary multi-level system and provide an illustration for a slightly more complicated three-level system, where the $(2,-1)$-contribution (DQC) can also exist outside of the cavity.
Discussion and conclusions.—We have presented an efficient and broadly applicable framework for nonlinear polariton spectroscopy, based on a mean-field evolution of the coupled light-matter system. The formalism extends conventional free-space spectroscopy by accounting for the feedback of the molecular polarization onto the cavity field and allows for an understanding of the polariton response beyond the stationary limit F. Ribeiro et al. (2018). Expanding the cavity field and molecular density matrix in terms of the input fields’ amplitudes and phases allows for a systematic understanding of the various contributions and excitation pathways to the nonlinear response.
The formalism offers potential for extension and generalization via multiple avenues. For instance, while our current treatment relies on a semiclassical description, therefore restricting it to a coherent state of light within the cavity, a valuable extension could involve capturing nonclassical photon statistics, e.g., by using cumulant expansion techniques Fowler-Wright et al. (2023). Other areas of interest are the study of spatially-resolved nonlinear polariton spectroscopy, which would involve generalizing the zero-dimensional cavity Eq. (3) to the 3D set of cavity modes afforded by classical electromagnetic simulations Taflove and Hagness (2005); Yang et al. (2023); Balasubrahmaniyam et al. (2023); Xu et al. (2023), or cavities exhibiting inherent nonlinearities Finkelstein-Shapiro et al. (2023). Furthermore, while we have mainly focused on pump-probe spectroscopy involving two pulses, the formalism can also be readily applied to 2D or multidimensional coherent spectroscopy by considering three or more pulses driving the cavity. Finally, let us emphasize that our approach is not limited to molecular polaritons, but could also be extended to other strongly coupled light-matter platforms, such as exciton-polaritons in inorganic semiconductor microcavities Fryett et al. (2018) or atomic gases.
Acknowledgments.—We are grateful to Wei Xiong for providing valuable feedback and comments on the manuscript. This research was primarily supported by the Air Force Office of Scientific Research (AFOSR) through the Multi-University Research Initiative (MURI) program no. FA9550-22-1-0317.

### v4

@@
 The formalism extends traditional free-space nonlinear spectroscopy by incorporating the feedback of matter onto the light field via the induced polarization.
-We demonstrate the utility of the framework by applying it to the calculation of pump-probe polariton spectra and show how, by storing the pulses, the cavity facilitates additional excitation pathways which are not possible in free space.
-Our method, which does not scale with $\mathcal{N}$, is broadly applicable and can be extended to a wide range of current experiments investigating the dynamical nonlinear response of hybrid light-matter states.
-pacs:
-42.50.Pq, 42.65.-k, 42.65.Sf
+We demonstrate the utility of the framework by applying it to the calculation of pump-probe polariton spectra and show how, by storing the pulses, the cavity facilitates additional excitation pathways, which can be used to isolate purely bright state contributions.
+Our method, which does not scale with $\mathcal{N}$, is broadly applicable and can be extended to model a wide range of current experiments investigating the dynamical nonlinear response of hybrid light-matter states.
 Figure 1: (a) Schematics.
@@
 (b) Key distinction between conventional, free-space nonlinear spectroscopy and nonlinear cavity spectroscopy, concerning the evolution of the molecular density matrix.
-For cavity spectroscopy, the feedback of the molecular polarization onto the cavity field leads to a coupling of the $n$-th order molecular density matrix $\vec{\rho}^{(n)}$ to all lower orders in the cavity field, as opposed to the free-space case, where the interaction is always linear in the laser field that drives the molecules.
-(c) Pump and probe pulses entering the cavity with time delay $\tau_{\Delta}$ get stored and prolonged in the cavity on a time scale given by the cavity lifetime $\kappa^{-1}$, therefore allowing the cavity field created by the probe ($\alpha^{(0)(1)}$) to act before the pump ($\alpha^{(1)(0)}$).
+For cavity spectroscopy, the feedback of the molecular polarization onto the cavity field leads to a coupling of the $n$th order molecular density matrix $\vec{\rho}^{(n)}$ to all lower orders in the cavity field, as opposed to the free-space case, where the interaction is always linear in the laser field that drives the molecules.
+(c) Pump and probe pulses entering the cavity at time delay $\tau_{\Delta}$ get stored and prolonged in the cavity on a time scale given by the cavity lifetime $\kappa^{-1}$, thereby allowing the (linear) cavity field created by the probe ($\alpha^{(0)(1)}$) to act before the pump ($\alpha^{(1)(0)}$).
 Additional phases $\Phi_{p,p^{\prime}}$ imprinted on the pulses can be utilized to separate excitation pathways in Liouville space.
@@
 (2022); Russo et al.
+(2024); Michail et al.
 (2024) and vibrational Xiang et al.
@@
 (2023) nonlinear polariton spectroscopy have provided new insight into the nonlinear response of molecular polaritons by revealing the dynamics of the complex interactions and energy relaxation pathways.
-However, significant challenges and open questions remain in understanding the nonlinear polariton response, particularly regarding the short-time dynamical behavior, i.e., before the cavity photons, and thus the polariton states, have decayed ($\sim$ fs in experiments).
+However, significant challenges and open questions remain in understanding the nonlinear polariton response, particularly regarding the short-time dynamical behavior, i.e., before the cavity photons, and thus the polariton states, have decayed ($\sim$ fs-ps in experiments).
 The development of an efficient formalism that can address these questions is crucial for advancing the understanding of polariton dynamics across different time scales as observed in pump-probe experiments.
-While the theoretical formalism of nonlinear spectroscopy involving the interaction between laser light and matter in free space is well-established Mukamel (1995); Yuen-Zhou et al.
+While the theoretical formalism of nonlinear spectroscopy involving the interaction between laser light and matter in free space is well-established Mukamel (1995); Boyd (2008); Yuen-Zhou et al.
 (2014); Jonas (2003); Cho (2008); Gelin et al.
-(2009), recent works have aimed at constructing a framework for the theory of of nonlinear electronic and vibrational polariton spectroscopy Saurabh and Mukamel (2016); F.
+(2009), recent works have aimed at constructing a framework for the theory of nonlinear electronic and vibrational polariton spectroscopy Saurabh and Mukamel (2016); F.
 Ribeiro et al.
@@
 Our approach is based on a mean-field evolution of the coupled light-matter system Fowler-Wright et al.
-(2022), related to semiclassical Maxwell-Liouville approaches which have been developed, e.g., in semiconductor optics and nanophotonics Jahnke et al.
+(2022); Fowler-Wright (2024), related to semiclassical Maxwell-Liouville approaches which have been developed, e.g., in semiconductor optics and nanophotonics Jahnke et al.
 (1996); Scully and Zubairy (1997); Lopata and Neuhauser (2009); Šindelka (2010); Sukharev and Nitzan (2011); Li et al.
 (2018); Chen et al.
-(2019).
+(2019); Bhat and Sipe (2001, 2006).
 We perturbatively expand both cavity field and density matrix in the input fields driving the cavity and show that the formalism can be regarded as an extension of the conventional free-space perturbative nonlinear spectroscopy framework by Mukamel Mukamel (1995), now incorporating nonlinear contributions to the field resulting from the feedback of the molecules onto the cavity dynamics.
-In addition, expansion of the nonlinear response in terms of phase components allows us to separate different excitation pathways in Liouville space and identify new pathways enabled by the cavity, using the widely-adopted toolkit of double-sided Feynman diagrams.
+Expansion of the nonlinear response in terms of phase components allows us to separate different excitation pathways in Liouville space and identify purely bright state contributions, using the widely-adopted toolkit of double-sided Feynman diagrams.
 Mean-field approach.—We consider an ensemble of $\mathcal{N}$ molecules collectively coupled to the single confined mode of an optical cavity at frequency $\omega_{c}$ [see schematics in Fig.
@@
 Each molecule $j$ is described by an (in principle arbitrary) Hamiltonian $\mathcal{H}_{0}^{j}$ with a corresponding dipole operator $\hat{\mu}_{j}$.
-The interaction between molecules and the cavity field is described by a Dicke-type Hamiltonian such that the total Hamiltonian describing the full quantum evolution of the coupled light-matter system expresses as (we set $\hbar\equiv 1$)
+The interaction between the molecules and the cavity field is described by a Dicke-type Hamiltonian such that the total Hamiltonian describing the full quantum evolution of the coupled light-matter system expresses as (we set $\hbar\equiv 1$)
 $\mathcal{H}=\omega_{c}\hat{a}^{\dagger}\hat{a}+\sum_{j=1}^{\mathcal{N}}\mathcal{H}_{0}^{j}+E_{0}(\hat{a}+\hat{a}^{\dagger})\sum_{j=1}^{\mathcal{N}}\hat{\mu}_{j},$ (1)
@@
 (2022).
-Additionally, for illustration, we take all molecules to be identical, i.e., $\hat{\mu}_{j}=\hat{\mu}$, although extension to disordered ensembles offers no conceptual difficulty.
+Additionally, for illustration, we take all molecules to be identical, i.e., $\mathcal{H}_{0}^{j}=\mathcal{H}_{0}$, $\hat{\mu}_{j}=\hat{\mu}$, although extension to disordered ensembles offers no conceptual difficulty.
 The simplified mean-field Hamiltonian then expresses as
 $\mathcal{H}_{\text{MF}}=\mathcal{H}_{0}+E_{0}\left(\alpha(t)+\alpha^{*}(t)\right)\hat{\mu}.$ (2)
-The dynamical evolution of the (now single-particle) molecular density matrix $\rho$ is then governed by the master equation $\partial_{t}\rho=\mathrm{i}[\rho,\mathcal{H}_{\text{MF}}]+\mathcal{D}[\rho]$.
-Here, the dissipator $\mathcal{D}[\rho]$ represents possible additional loss channels of the molecule (e.g., dephasing at rate $\gamma_{\phi}$, non-radiative decay,$\ldots$) that may be added, e.g., in Lindblad form.
+The dynamical evolution of the (now single-molecule) density matrix $\rho$ is governed by the master equation $\partial_{t}\rho=\mathrm{i}[\rho,\mathcal{H}_{\text{MF}}]+\mathcal{D}[\rho]$.
+Here, the dissipator $\mathcal{D}[\rho]$ represents possible additional loss channels of the molecule (e.g., dephasing at rate $\gamma_{\phi}$, non-radiative decay,…) that may be added, e.g., in Lindblad form.
 However, a complete description of the mean-field problem also involves solving the classical equation for the cavity field amplitude which undergoes damping at rate $\kappa$ (photon loss)
 $\dot{\alpha}(t)$ $=-\left(\frac{\kappa}{2}+\mathrm{i}\omega_{c}\right)\alpha(t)-\mathrm{i}\mathcal{N}E_{0}P(t)-\eta f(t)\mathrm{e}^{-\mathrm{i}\omega_{\ell}t},$ (3)
-containing the induced molecular polarization $P(t)=\mathrm{Tr}[\hat{\mu}\rho(t)]$ created by all $\mathcal{N}$ molecules.
+containing the induced molecular polarization $P(t)=\mathrm{Tr}[\hat{\mu}\rho(t)]$ of all $\mathcal{N}$ molecules.
 Altogether, this describes a self-consistent, semiclassical evolution of the coupled light-matter system, where the matter component evolves quantum mechanically, while the part describing the electromagnetic field undergoes classical evolution and is driven by both the input field as well as by the collective molecular polarization feeding back into the cavity field.
+While Eq.
+(3) assumes a simplified, single-mode description of the electromagnetic field, we discuss the straightforward extension to multimode cavities in the Supplementary Material (SM) sm .
 Perturbative expansion.—While the above set of mean-field equations can be readily solved numerically, there is generally no closed analytical solution due to the nonlinearity of the molecular transitions.
-To gain more insight into the structure of the nonlinear response, we therefore proceed in analogy to standard nonlinear spectroscopy with a perturbative expansion of both the field and the matter density matrix in terms of the input field amplitude, i.e.,
+To gain more insight into the structure of the nonlinear response, we therefore proceed, akin to standard nonlinear spectroscopy, with a perturbative expansion of both cavity field and matter density matrix in terms of the input field amplitude, i.e.,
 $\alpha(t)=\sum_{n=0}^{\infty}\eta^{n}\alpha^{(n)}(t),\quad\rho(t)=\sum_{n=0}^{\infty}\eta^{n}\rho^{(n)}(t).$ (4)
-From this, also other perturbative quantities can be derived such as, e.g., the $n$-th order polarization $P^{(n)}(t)=\mathrm{Tr}[\hat{\mu}\rho^{(n)}(t)]$  [^1] Note that the $\rho^{(n)}$ are not true density matrices, i.e., they do not obey the properties of density matrices such as the conservation of the trace.
-Furthermore, for both numerical and analytical considerations, it is advantageous to convert the master equation for the density matrix onto matrix-vector form, i.e., map the density matrix $\rho$ onto a vector $\vec{\rho}$ and the superoperator acting on the density matrix onto a matrix (also referred to as Liouville space) [see Supplementary Material (SM) sm for details on vectorization] Mukamel (1995); Horn and Johnson (1994); Am-Shallem et al.
+From this, also other perturbative quantities can be derived such as, e.g., the $n$th order polarization $P^{(n)}(t)=\mathrm{Tr}[\hat{\mu}\rho^{(n)}(t)]$  [^1] Note that the $\rho^{(n)}$ are not true density matrices, i.e., they do not obey the properties of density matrices such as the conservation of the trace.
+Furthermore, for both numerical and analytical considerations, it is advantageous to convert the master equation for the density matrix onto matrix-vector form, i.e., map the density matrix $\rho$ onto a vector $\vec{\rho}$ and the superoperator acting on the density matrix onto a matrix (also referred to as Liouville space) [see SM sm for details on vectorization] Mukamel (1995); Horn and Johnson (1994); Am-Shallem et al.
 (2015).
-Then, one obtains a closed system for the cavity field amplitude and the vectorized density matrix up to $n$-th order
+Then, one obtains a closed system for the cavity field amplitude and the vectorized density matrix up to $n$th order
 $\dot{\alpha}^{(n)}$ $=-\left(\frac{\kappa}{2}+\mathrm{i}\omega_{c}\right)\alpha^{(n)}-\mathrm{i}\mathcal{N}E_{0}P^{(n)}-\delta_{1n}f(t)\mathrm{e}^{-\mathrm{i}\omega_{\ell}t},$ (5a)
 $\dot{\vec{\rho}}^{(n)}$ $=-\mathrm{i}\mathcal{L}_{0}\vec{\rho}^{(n)}-\mathrm{i}E_{0}\sum_{j=0}^{n}\left(\alpha^{(n-j)}+\alpha^{(n-j)*}\right)\mathcal{L}_{\mu}\vec{\rho}^{(j)},$ (5b)
-in terms of the vectorized commutators $[\mathcal{H}_{0},\rho^{(n)}]\to\mathcal{L}_{0}\vec{\rho}^{(n)}$ and $[\hat{\mu},\rho^{(j)}]\to\mathcal{L}_{\mu}\vec{\rho}^{(j)}$ where $\mathcal{L}_{0}$, $\mathcal{L}_{\mu}$ are the free molecular and interaction Liouvillians, respectively.
+in terms of the vectorized commutators $[\mathcal{H}_{0},\rho^{(n)}]\to\mathcal{L}_{0}\vec{\rho}^{(n)}$ and $[\hat{\mu},\rho^{(j)}]\to\mathcal{L}_{\mu}\vec{\rho}^{(j)}$, where $\mathcal{L}_{0}$, $\mathcal{L}_{\mu}$ are the free molecular and interaction Liouvillians, respectively.
 In particular, Eqs.
-(5) indicate that only the first-order cavity field is driven directly by the input fields while for higher orders, the molecular density matrix is driven by lower orders of the cavity field.
+(5) indicate that only the first-order cavity field is driven directly by the input pulse while at higher orders, the molecular density matrix is driven by lower orders of the cavity field.
 Importantly, Eq.
@@
 In other words, the electric field experienced by the molecules is not the laser field interacting with the cavity from outside, but a distorted version of it due to the evolving material polarization inside the cavity.
-While for conventional spectroscopy with well-separated pulses, only a single multidimensional integral is obtained as solution for the $n$-th order density matrix, many additional integrals arise in the cavity scenario ($12^{n-1}-$ 1, see diagram in Fig.
-2 and SM sm ).
+While for conventional spectroscopy with well-separated pulses, only a single multidimensional integral is obtained as solution for the $n$th order density matrix, many additional integrals arise in the cavity scenario ($12^{n-1}-$ 1, see diagram in Fig.
+2).
 Also, a straightforward analysis of Eqs.
@@
 Figure 2: Density matrix evolution.
-Graph illustrating the possible integrals contributing to the dynamical evolution of the $n$-th order density matrix $\rho^{(n)}(t)$.
+Graph illustrating the possible integrals contributing to the dynamical evolution of the $n$th order molecular density matrix $\rho^{(n)}(t)$.
 The orange path on the right illustrates the integral that is obtained for conventional free-space spectroscopy, corresponding to a sequence of first-order interactions linear in the input.
@@
 Figure 3: Differential transmission.
-DT spectra $\Delta T(\omega)$ for a 2LS with resonance frequency $\omega_{0}$ as a function of the delay time between the pump and probe pulse $\tau_{\Delta}$ for (a) weak ($\gamma_{\phi}=0$), and (b) strong ($\gamma_{\phi}=\kappa$) dephasing of the electronic transition.
+DT spectra $\Delta T(\omega)$ for a 2LS with resonance frequency $\omega_{0}$ as a function of the delay time between the pump and probe pulse $\tau_{\Delta}$ for (a) weak ($\gamma_{\phi}=0$), and (b) strong ($\gamma_{\phi}=\kappa$) dephasing of the transition.
 The green dashed lines show the location of the (linear response) polariton frequencies.
 The black curve in (b) shows a cross section through the DT at $\tau_{\Delta}=4\kappa^{-1}$.
-The plots in (c) and (d) show the dynamics of the population and polarization $|P(t)|^{2}$ created by the initial pump pulse arriving at $\tau_{p}=2\kappa^{-1}$ which are then probed by the second pulse after $\tau_{\Delta}$.
-The 2LS is initialized in the ground state, with parameters $g\sqrt{\mathcal{N}}=3\kappa$, $\tau_{w}=0.05\kappa^{-1}$, $\omega_{0}=\omega_{c}$, and central frequency $\omega_{p}=\omega_{p^{\prime}}=\omega_{0}$ for both pulses.
+The results are plotted in a frame rotating at the central pulse frequency.
+Panels (c) and (d) depict the dynamics of the population and polarization $|P(t)|^{2}$ created by the initial pump pulse arriving at $\tau_{p}=2\kappa^{-1}$ which are then probed by the second pulse at delay $\tau_{\Delta}$.
+The 2LS is initialized in the ground state, with parameters $g\sqrt{\mathcal{N}}=3\kappa$, $\tau_{w}=0.05\kappa^{-1}$, $\omega_{0}=\omega_{c}$, and carrier frequency $\omega_{p}=\omega_{p^{\prime}}=\omega_{0}$ for both pulses.
 Differential transmission.—We now consider two pulses driving the cavity (pump $p$ and probe $p^{\prime}$), with input fields proportional to $\eta_{j}f_{j}(t-\tau_{j})\mathrm{e}^{-\mathrm{i}\omega_{j}t}$ where $\eta_{j}$, $f_{j}$, $\omega_{j}$, $\tau_{j}$ for $j\in\{p,p^{\prime}\}$ denote the amplitude, envelope, carrier frequency, and arrival time of the pump and probe pulses, respectively.
-We assume $\tau_{p^{\prime}}>\tau_{p}$, with a time delay between pump and probe of $\tau_{\Delta}=\tau_{p^{\prime}}-\tau_{p}$ [see sketch in Fig.
+We assume a time delay between pump and probe of $\tau_{\Delta}=\tau_{p^{\prime}}-\tau_{p}\geq 0$ [see sketch in Fig.
 1(c)].
@@
 $\rho(t)$ $=\sum_{n,m=0}^{\infty}\eta_{p}^{n}\eta_{p^{\prime}}^{m}\rho^{(n)(m)}(t),$ (6b)
-where the first and second indeces refer to the orders in the pump and probe fields, respectively.
+where the first and second indices refer to the orders in the pump and probe fields, respectively.
 This allows us to derive a perturbative set of equations for $\alpha^{(n)(m)}$, $\vec{\rho}^{(n)(m)}$ similar to Eqs.
@@
 (a), (b) Plot of phase contributions $(0,1)$ and $(2,-1)$ to $\Delta T(\omega)$ as a function of the delay time between the pulses.
-The $(2,-1)$-contribution decays on a time scale of $\kappa^{-1}$, as it relies on the probe acting prior to the pump.
+The $(2,-1)$-contribution decays on a time scale of $\sim(\kappa+\gamma_{\phi})^{-1}$, as it relies on the probe acting on the coherence created by the pump, therefore probing the bright state only.
 The green dashed lines show the (linear response) polariton frequencies.
 We chose $g\sqrt{\mathcal{N}}=3\kappa$, $\gamma_{\phi}=0.1\kappa$, $\omega_{p}=\omega_{p^{\prime}}=\omega_{0}$, and a temporal width of $\tau_{w}=0.1\kappa^{-1}$ for both pulses.
-The double-sided Feynman diagrams in (c) and (d) represent processes contributing to the different phase combinations of $\alpha^{(2)(1)}$.
+The double-sided Feynman diagrams in (c) and (d) represent processes contributing to the different phase combinations of the third-order cavity field $\alpha^{(2)(1)}$.
 The dotted arrows indicate the decay of the coherence.
-From the solution for the total cavity field, the cavity transmission and reflection can be obtained using standard input-output theory Steck (2007).
+From the solution for the total cavity field, the cavity transmission and reflection can be obtained using standard input-output theory Steck (2007) (valid as long as the system is not in the ultrastrong coupling regime Ciuti and Carusotto (2006); Frisk Kockum et al.
+(2019)).
 Since nonlinearities are typically weak, instead of absolute quantities, differential (nonlinear) corrections to the response are commonly considered.
-For the collinear geometry in question, we define the differential transmission (DT) spectrum by subtracting the (normalized) probe transmission without the pump from the probe transmission in the presence of the pump pulse
+For the collinear geometry in question, we define the differential transmission (DT) spectrum by subtracting the (normalized) probe transmission without the pump from the probe transmission in the presence of the pump field
 $\Delta T(\omega)=T_{p^{\prime}}^{\text{pump\, on}}(\omega)-T_{p^{\prime}}^{\text{pump\, off}}(\omega),$ (7)
-where the transmission is related to the intracavity field $T_{p^{\prime}}=|(\kappa/2)\alpha(\omega)/[\eta_{p^{\prime}}f_{p^{\prime}}(\omega)]|^{2}$.
+where the transmission is related to the intracavity field as $T_{p^{\prime}}(\omega)=|(\kappa/2)\alpha(\omega)/[\eta_{p^{\prime}}f_{p^{\prime}}(\omega)]|^{2}$.
 The contribution $T_{p^{\prime}}^{\text{pump\, on}}(\omega)$ depends on the time delay between the pulses $\tau_{\Delta}$.
@@
 $\Delta T(\omega)\approx\left(\frac{\kappa}{2}\right)^{2}\eta_{p}^{2}\frac{2\mathrm{Re}\left[\alpha^{(0)(1)*}(\omega)\alpha^{(2)(1)}(\omega)\right]}{f_{p^{\prime}}(\omega)^{2}},$ (8)
-where the key quantity is the third-order cavity field $\alpha^{(2)(1)}$.
+describing interference [four-wave mixing (4WM)] between the probe (acting as local oscillator) and the third-order cavity field $\alpha^{(2)(1)}$.
 Beware that Eq.
-(8) is not the usual pump-probe signal; instead, it is a signal which contains different phase contributions (see paragraph below).
+(8) is not the usual pump-probe signal; instead, it contains different phase contributions (see paragraph below).
 The DT is plotted as a function of delay time in Figs.
-3(a), (b) for two-level systems (2LSs) with transition dipole moment $\mu_{ge}$ coupled to the cavity at strength $g=\mu_{ge}E_{0}$, for various degrees of dephasing of the transition.
+3(a), (b) for two-level systems (2LSs) with transition dipole moment $\mu_{ge}$ coupled to the cavity at strength $g=\mu_{ge}E_{0}$, for two different degrees of dephasing.
 The plots below in Figs.
@@
 At early times, oscillatory behavior around the polariton frequencies is observed, corresponding to coherences between upper and lower polariton.
-Dephasing leads to a decay of coherence and the creation of stationary population in the molecules, even after the cavity photon has decayed.
+Dephasing leads to a faster decay of coherence and the creation of stationary population in the molecules, even after the cavity photon has decayed.
 In the long-time limit, this manifests as a contracted Rabi splitting, characterized by a DT as shown by the black cross-section curve in Fig.
@@
 (2018).
-Separating excitation pathways.—The nonlinear signal stemming from the $n$-th order density matrix generally contains many possible excitation pathways.
+Separating excitation pathways.—The nonlinear signal stemming from the $n$th order density matrix generally contains many possible excitation pathways.
 For instance, even for free-space spectroscopy of a 2LS, there are eight Feynman diagrams contributing to the third-order nonlinear response Mukamel (1995).
-In experiments, it is therefore highly desired to isolate certain excitation pathways and thereby understand the different contributions to the nonlinear signal, which, owing to the collinear geometry in question, can be achieved by phase cycling.
-This method is equivalent to the more standard phase matching procedure, which, however obfuscates our formalism as it requires the description of a multimode cavity Yuen-Zhou et al.
+In experiments, it is therefore highly desired to isolate certain excitation pathways and thereby understand the different contributions to the nonlinear signal, which here, owing to the collinear geometry in question, can be achieved by phase cycling.
+This method is equivalent to the more standard phase matching procedure with oblique pulses, which, however obfuscates our formalism as it requires the description of a multimode cavity Yuen-Zhou et al.
 (2014); Jonas (2003); Hybl et al.
@@
 $\Delta T_{\vec{v}}(\omega)\approx\left(\frac{\kappa}{2}\right)^{2}\eta_{p}^{2}\frac{2\mathrm{Re}\left[\alpha^{(0)(1)*}_{(0,-1)}(\omega)\alpha^{(2)(1)}_{\vec{v}}(\omega)\right]}{f_{p^{\prime}}(\omega)^{2}},$ (10)
-with the total signal simply being obtained as the sum of all contributions $\Delta T=\sum_{\vec{v}}\Delta T_{\vec{v}}$.
-One finds that there are two possible phase combinations contributing to $\alpha^{(2)(1)}$: $(0,1)$, and $(2,-1)$ [and the complex conjugate processes $(0,-1)$, $(-2,1)$], which are shown in Figs.
-4(a), (b), and are associated with pump-probe and double-quantum coherence (DQC) spectroscopies, respectively.
-The first contribution [$(0,1)\equiv\pm\Phi_{p}\mp\Phi_{p}+\Phi_{p^{\prime}}$] is the ‘standard’ pump-probe, where populations and zeroth-order quantum coherences are created by twofold application of the pump with opposite phase [see Feynman diagrams in Fig.
+where $(0,-1)$ is the only phase contribution for the conjugate probe field sm and the total signal is obtained as $\Delta T=\sum_{\vec{v}}\Delta T_{\vec{v}}$.
+Two phase combinations contribute to $\alpha^{(2)(1)}$: $(0,1)$ and $(2,-1)$ [along with the complex conjugates $(0,-1)$ and $(-2,1)$], corresponding to pump–probe and double-quantum coherence (DQC) spectroscopies, respectively [Figs.
+4(a,b)].
+The pump–probe pathway $(0,1)\equiv\pm\Phi_{p}\!\mp\!\Phi_{p}\!+\!\Phi_{p^{\prime}}$ creates populations and zeroth-order quantum coherences via two pump pulses of opposite phase [Fig.
 4(c)].
-The coherent superposition between upper and lower polaritons then undergo Rabi oscillations which are detected by the probe.
-The second phase combination [($(2,-1)\equiv 2\Phi_{p}-\Phi_{p^{\prime}}$] is negligible outside of the cavity, as it implies that the 2LS is excited twice.
-However, inside the cavity, since the pulses get extended on a time scale of $\kappa^{-1}$ [cf.
-Fig.
-1(c)], the probe can act before the pump, therefore enabling this phase combination to arise within $\kappa^{-1}$.
-In the SM sm , we derive the phase contributions up to third order for an arbitrary multi-level system and provide an illustration for a slightly more complicated three-level system, where the $(2,-1)$-contribution (DQC) can also exist outside of the cavity.
+The DQC pathway $(2,-1)\equiv 2\Phi_{p}\!-\!\Phi_{p^{\prime}}$ is negligible outside the cavity because it requires the two-level system to be excited twice.
+However, inside the cavity, the probe can act on the pump-induced coherence within the lifetime of the bright state $(\kappa+\gamma_{\phi})^{-1}$ [left diagram in Fig.
+4(d)], enabling this pathway to arise within this time interval.
+The right diagram in Fig.
+4(d) does not contribute to the probe DT since the pump has not acted yet.
+Consequently, the $(2,-1)$ pathway isolates the bright state only, in contrast to the $(0,1)$ pathway, which accesses both bright and dark states (and probes only dephased dark states in the long time limit).
+Eq.
+(10) satisfies the phase matching conditions for degenerate 4WM with two pulses Yuen-Zhou et al.
+(2014).
 Discussion and conclusions.—We have presented an efficient and broadly applicable framework for nonlinear polariton spectroscopy, based on a mean-field evolution of the coupled light-matter system.
@@
 (2018).
-Expanding the cavity field and molecular density matrix in terms of the input fields’ amplitudes and phases allows for a systematic understanding of the various contributions and excitation pathways to the nonlinear response.
+Expanding the cavity field and molecular density matrix in terms of the input fields’ amplitudes and phases provides a systematic understanding of the various contribution to the nonlinear polariton response, and allows to separate purely bright state contributions from dark reservoir states.
 The formalism offers potential for extension and generalization via multiple avenues.
-For instance, while our current treatment relies on a semiclassical description, therefore restricting it to a coherent state of light within the cavity, a valuable extension could involve capturing nonclassical photon statistics, e.g., by using cumulant expansion techniques Fowler-Wright et al.
-(2023).
+For instance, while our current treatment relies on a semiclassical mean-field description, valuable extensions could involve capturing nonclassical photon statistics or many-body correlations among particles, e.g., by using cumulant or cluster expansion techniques Fowler-Wright et al.
+(2023); Axt and Stahl (1994).
 Other areas of interest are the study of spatially-resolved nonlinear polariton spectroscopy, which would involve generalizing the zero-dimensional cavity Eq.
-(3) to the 3D set of cavity modes afforded by classical electromagnetic simulations Taflove and Hagness (2005); Yang et al.
+(3) to the 3D set of cavity modes afforded by classical electromagnetic simulations Gruetzmacher and Scherer (2003); Zhou et al.
+(2024); Yang et al.
 (2023); Balasubrahmaniyam et al.
@@
 (2023).
-Furthermore, while we have mainly focused on pump-probe spectroscopy involving two pulses, the formalism can also be readily applied to 2D or multidimensional coherent spectroscopy by considering three or more pulses driving the cavity.
+Furthermore, while we have focused on pump-probe spectroscopy involving two pulses, the formalism can be readily applied to 2D or multidimensional coherent spectroscopy, which may also help to disentangle pure electronic dephasing from vibronic relaxation channels.
 Finally, let us emphasize that our approach is not limited to molecular polaritons, but could also be extended to other strongly coupled light-matter platforms, such as exciton-polaritons in inorganic semiconductor microcavities Fryett et al.
-(2018) or atomic gases.
+(2018) or atomic ensembles.
 Acknowledgments.—We are grateful to Wei Xiong for providing valuable feedback and comments on the manuscript.
@@
 FA9550-22-1-0317.
+Supplemental Material to Nonlinear semiclassical spectroscopy of ultrafast molecular polariton dynamics
+S1 Vectorization/mapping to Liouville space
+It is advantageous to convert the von-Neumann (master) equation for the density matrix onto matrix-vector form, i.e., map the $n\times n$ density matrix $\rho$ onto an $n^{2}$ vector $\vec{\rho}$ and the superoperator acting on the density matrix onto an $n^{2}\times n^{2}$ matrix $\mathcal{L}$ (referred to as Liouville space).
+For instance, for a two-level system, the density matrix is mapped onto a column vector as
+$\rho=\begin{pmatrix}\rho_{11}&\rho_{12}\\ \rho_{21}&\rho_{22}\end{pmatrix}\quad\to\quad\vec{\rho}\equiv|\rho\rangle\rangle=\begin{pmatrix}\rho_{11}\\ \rho_{12}\\ \rho_{21}\\ \rho_{22}\end{pmatrix}.$ (S1)
+This can be done in a more formal way in a procedure known in the literature as vec-ing Am-Shallem et al. (2015); Horn and Johnson (1994), which has the following rules
+  1. 1.
+    A left multiplication of the matrix $\rho$ by an $n\times n$ matrix $A$, i.e., $A\rho$, is equivalent to an operation on the vector $\vec{\rho}$ by the $n^{2}\times n^{2}$ matrix $A\otimes I$, where $I$ is the $n\times n$ identity matrix, and $\otimes$ is the Kronecker product.
+  2. 2.
+    Similarly, a right multiplication of the matrix $\rho$ by an $n\times n$ matrix $B$, i.e., $\rho B$, is equivalent to an operation on the vector $\vec{\rho}$ by the $n^{2}\times n^{2}$ matrix $I\otimes B^{\top}$.
+  3. 3.
+    Finally, a combination of left and right matrices multiplication, $A\rho B$, is equivalent to an operation on the vector $\vec{\rho}$ by the $n^{2}\times n^{2}$ matrix $A\otimes B^{\top}$.
+With this, one can express general commutators as
+$[\mathcal{H},\rho]\to\left(\mathcal{H}\otimes I-I\otimes\mathcal{H}^{\top}\right)\vec{\rho}\equiv\mathcal{L}\vec{\rho}.$ (S2)
+S2 Cavity input-output relations
+Let us consider a (classical) input field $\expectationvalue{\hat{b}_{\mathrm{in},L}(t)}$ driving the cavity through the left mirror described by the Hamiltonian
+$\mathcal{H}_{d}=-\mathrm{i}\hbar\sqrt{\kappa_{L}}\expectationvalue{\hat{b}_{\mathrm{in},L}(t)}\hat{a}^{\dagger}+\mathrm{H.c.},$ (S3)
+where $\kappa_{L}$ describes the loss rate through the left mirror.
+For our purposes, we consider an input pulse described by $\expectationvalue{\hat{b}_{\mathrm{in},L}(t)}=\eta f(t)/\sqrt{\kappa_{L}}$ (in a frame rotating at the pulse frequency).
+The parameter $\eta$ can be related to the incident laser power $\mathcal{P}_{\text{in}}$ as $\eta=\sqrt{\mathcal{P}_{\text{in}}\kappa_{L}/(\hbar\omega_{\ell})}$, where $\omega_{\ell}$ is the pulse carrier frequency.
+Then, the cavity input-output relations written separately at the left ($L$) and right ($R$) mirror, respectively, are given by Steck (2007)
+$\hat{b}_{\mathrm{out},L}(t)-\hat{b}_{\mathrm{in},L}(t)$ $=\sqrt{\kappa_{L}}\hat{a},$ (S4a)
+$\hat{b}_{\mathrm{out},R}(t)-\hat{b}_{\mathrm{in},R}(t)$ $=\sqrt{\kappa_{R}}\hat{a},$ (S4b)
+from which the transmitted intensity can be straightforwardly derived as
+$T=\left|\frac{\expectationvalue{\hat{b}_{\mathrm{out},R}(t)}}{\expectationvalue{\hat{b}_{\mathrm{in},L}(t)}}\right|^{2}=\frac{\kappa_{R}\kappa_{L}}{\eta^{2}f(t)^{2}}|\alpha|^{2}.$ (S5)
+For a balanced cavity $\kappa_{L}=\kappa_{R}=\kappa/2$, this becomes
+$T=\frac{(\kappa/2)^{2}}{\eta^{2}f(t)^{2}}|\alpha|^{2},$ (S6)
+which holds in both linear and nonlinear regimes, as well as in the Fourier domain.
+It is important to note, however, that in the ultrastrong coupling regime, the input-output relations are significantly modified due to the breakdown of the rotating wave approximation (RWA).
+In this regime, where the collective coupling strength $g\sqrt{\mathcal{N}}$ becomes comparable to the cavity or matter transition frequencies, non-trivial corrections to the input-output relations must be incorporated that account for counter-rotating terms and the hybrid eigenmodes of the light-matter system Ciuti and Carusotto (2006); Frisk Kockum et al.
+(2019).
+S3 General form of pump-probe equations
+For a cavity driven by two pulses, the $(n)(m)$-th order equations for the field and molecular density matrix are given by (assuming initial conditions $\alpha(0)=\alpha^{(0)(0)}=0$)
+$\dot{\alpha}^{(n)(m)}$ $=-\left(\frac{\kappa}{2}+\mathrm{i}\omega_{c}\right)\alpha^{(n)(m)}-\mathrm{i}E_{0}\mathcal{N}\mathrm{Tr}[\hat{\mu}\rho^{(n)(m)}]-\delta_{1n}\delta_{0m}f_{p}(t-\tau_{p})\mathrm{e}^{-\mathrm{i}\omega_{p}t}-\delta_{0n}\delta_{1m}f_{p^{\prime}}(t-\tau_{p^{\prime}})\mathrm{e}^{-\mathrm{i}\omega_{p^{\prime}}t},$ (S7a)
+$\dot{\rho}^{(n)(m)}(t)$ $=-\mathrm{i}[\mathcal{H}_{0},\rho^{(n)(m)}]-\mathrm{i}E_{0}\sum_{j=0}^{n-1}\sum_{j^{\prime}=0}^{m-1}\left(\alpha^{(n-j)(m-j^{\prime})}+\alpha^{(n-j)(m-j^{\prime})*}\right)\left[\hat{\mu},\rho^{(j)(j^{\prime})}\right],$ (S7b)
+where we will generally assume identical pulse envelopes $f_{j}(t-\tau_{j})=\mathrm{e}^{-(t-\tau_{j})^{2}/\tau_{w}^{2}}$.
+In Liouville space, the equation for the density vector can then be expressed in compact form as
+$\dot{\vec{\rho}}^{(n)(m)}(t)=-\mathrm{i}\mathcal{L}_{0}\vec{\rho}^{(n)(m)}(t)-\mathrm{i}\sum_{j=0}^{n-1}\sum_{j^{\prime}=0}^{m-1}\mathcal{L}_{\text{int}}^{(n-j)(m-j^{\prime})}(t)\vec{\rho}^{(j)(j^{\prime})}(t),$ (S8)
+where we defined the interaction Liouvillian
+$\mathcal{L}_{\text{int}}^{(n-j)(m-j^{\prime})}(t)=E_{0}[\alpha^{(n-j)(m-j^{\prime})}(t)+\mathrm{c.c.}]\mathcal{L}_{\mu},$ (S9)
+and the vectorized commutators are given by
+$[\mathcal{H}_{0},\rho]\to\mathcal{L}_{0}\vec{\rho}$ $=[\mathcal{H}_{0}\otimes I-I\otimes\mathcal{H}_{0}^{\top}]\vec{\rho},$ (S10a)
+$[\hat{\mu},\rho]\to\mathcal{L}_{\mu}\vec{\rho}$ $=[\hat{\mu}\otimes I-I\otimes\hat{\mu}^{\top}]\vec{\rho}.$ (S10b)
+S4 Differential transmission
+We define the differential transmission (DT) as the difference in the probe transmission with and without the presence of the pump pulse
+$\Delta T(\omega)=T_{p^{\prime}}^{\text{pump\, on}}(\omega)-T_{p^{\prime}}^{\text{pump\, off}}(\omega).$ (S11)
+Considering only the lowest (third) order correction of the probe field due to the pump, the normalized transmitted probe intensities can be expressed from input-output relations in Fourier domain (see section S2) as
+$T_{p^{\prime}}^{\text{pump\, off}}(\omega)=\left(\frac{\kappa}{2}\right)^{2}\frac{|\alpha^{(0)(1)}(\omega)|^{2}}{f_{p^{\prime}}(\omega)^{2}},\qquad T_{p^{\prime}}^{\text{pump\, on}}(\omega)\approx\left(\frac{\kappa}{2}\right)^{2}\frac{|\alpha^{(0)(1)}(\omega)+\eta_{p}^{2}\alpha^{(2)(1)}(\omega)|^{2}}{f_{p^{\prime}}(\omega)^{2}}.$ (S12)
+From this, we can approximate the DT as
+$\Delta T(\omega)\approx\left(\frac{\kappa}{2}\right)^{2}\eta_{p}^{2}\frac{2\mathrm{Re}\left[\alpha^{(0)(1)*}(\omega)\alpha^{(2)(1)}(\omega)\right]}{f_{p^{\prime}}(\omega)^{2}}.$ (S13)
+This quantity is not the usual pump-probe signal but contains different phase contributions (see main text, paragraph ‘Differential transmission’).
+S5 Analytical solution of pump-probe dynamics for a 2LS
+For a two-level system, let us consider the $(2)(1)$-order equations for the cavity field and the density matrix (second order in pump, first in probe), required for the computation of the DT derived in Sec.
+S4 (assuming a matter initial state $\rho(0)=\ket{g}\bra{g}$ |)
+$\dot{\alpha}^{(2)(1)}(t)$ $=-\left(\frac{\kappa}{2}+\mathrm{i}\omega_{c}\right)\alpha^{(2)(1)}-\mathrm{i}g\mathcal{N}\rho_{21}^{(2)(1)},$ (S14a)
+$\dot{\rho}^{(2)(1)}_{21}(t)$ $=-\left(\frac{\gamma+\gamma_{\phi}}{2}+\mathrm{i}\omega_{0}\right)\rho_{21}^{(2)(1)}-\mathrm{i}g\alpha^{(2)(1)}+\mathrm{i}g\alpha^{(0)(1)}(\rho_{22}^{(2)(0)}-\rho_{11}^{(2)(0)})$ (S14b)
+$+\mathrm{i}g\alpha^{(1)(0)}(\rho_{22}^{(1)(1)}-\rho_{11}^{(1)(1)}),$
+where we considered dissipation affecting the molecules in the form of dephasing ($\gamma_{\phi}$) and spontaneous emission ($\gamma$).
+The third-order coherence $\rho_{21}^{(2)(1)}$ is driven by the population created by the pump $(2)(0)$ as well as by the population created by both pump and probe $(1)(1)$.
+We rewrite the equation for the third-order coherence as
+$\dot{\rho}^{(2)(1)}_{21,\tau}(t)=-\left(\frac{\gamma+\gamma_{\phi}}{2}+\mathrm{i}\omega_{0}\right)\rho_{21,\tau}^{(2)(1)}-\mathrm{i}g\alpha^{(2)(1)}_{\tau}+F_{\tau}(t),$ (S15)
+where we condensed the input affecting the third-order coherence into a single term $F_{\tau}(t)$ and by the index $\tau$ we indicate that the quantities depend on the delay time between the pulses $\tau_{\Delta}$.
+In Fourier space, we can therefore obtain the solution for $\alpha^{(2)(1)}_{\tau}(\omega)$
+$\dot{\alpha}^{(2)(1)}_{\tau}(\omega)=\frac{-\mathrm{i}g\mathcal{N}F_{\tau}(\omega)}{\left[\mathrm{i}(\omega_{c}-\omega)+\frac{\kappa}{2}\right]\left[\mathrm{i}(\omega_{0}-\omega)+\frac{\gamma+\gamma_{\Phi}}{2}\right]+g^{2}\mathcal{N}}.$ (S16)
+From the equations of motion, we can relate second-order quantities to first-order quantities in frequency space as ($\ast$ denotes the convolution operation)
+$\rho_{22}^{(2)(0)}(\omega)-\rho_{11}^{(2)(0)}(\omega)$ $=\frac{4\mathrm{i}g}{\omega+\mathrm{i}\gamma}\mathrm{Im}\left[\left(\alpha^{(1)(0)}\ast\rho_{12}^{(1)(0)}\right)(\omega)\right],$ (S17a)
+$\rho_{22}^{(1)(1)}(\omega)-\rho_{11}^{(1)(1)}(\omega)$ $=\frac{4\mathrm{i}g}{\omega+\mathrm{i}\gamma}\left\{\mathrm{Im}\left[\left(\alpha^{(0)(1)}\ast\rho_{12}^{(1)(0)}\right)(\omega)\right]+\mathrm{Im}\left[\left(\alpha^{(1)(0)}\ast\rho_{12}^{(0)(1)}\right)(\omega)\right]\right\},$ (S17b)
+and therefore, we can express $F_{\tau}(\omega)$ as a sum of double convolutions
+$F_{\tau}(\omega)=$ $-\frac{4g^{2}}{\omega+\mathrm{i}\gamma}\Bigl{\{}\left[\alpha^{(0)(1)}\ast\mathrm{Im}\left[\alpha^{(1)(0)}\ast\rho_{12}^{(1)(0)}\right]\right]_{\tau}(\omega)\!+\!\left[\alpha^{(1)(0)}\ast\mathrm{Im}\left[\alpha^{(0)(1)}\ast\rho_{12}^{(1)(0)}\right]\right]_{\tau}(\omega)\!$
+$+\!\left[\alpha^{(1)(0)}\ast\mathrm{Im}\left[\alpha^{(1)(0)}\ast\rho_{12}^{(0)(1)}\right]\right]_{\tau}(\omega)\Bigr{\}},$ (S18)
+in terms of the first-order quantities in the pump field
+$\alpha^{(1)(0)}(\omega)=\frac{-\tilde{f}_{p}(\omega-\omega_{p})}{\left[\frac{\kappa}{2}+\mathrm{i}(\omega_{c}-\omega)\right]+\frac{g^{2}\mathcal{N}}{\frac{\gamma+\gamma_{\phi}}{2}+\mathrm{i}(\omega_{0}-\omega)}},\qquad\rho_{21}^{(1)(0)}(\omega)=\frac{-\mathrm{i}g\alpha^{(1)(0)}(\omega)}{\mathrm{i}(\omega_{0}-\omega)+\frac{\gamma+\gamma_{\phi}}{2}},$ (S19)
+and similarly for the terms linear in the probe. The Fourier transform of the Gaussian pulse shapes is given by
+$\tilde{f}_{j}(\omega)=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\infty}\mathrm{d}t\,\mathrm{e}^{-\frac{(t-\tau_{j})^{2}}{2\tau_{w}^{2}}}\mathrm{e}^{\mathrm{i}\omega t}=\tau_{w}\mathrm{e}^{-\omega^{2}\tau_{w}^{2}/2}\mathrm{e}^{\mathrm{i}\omega\tau_{j}},$ (S20)
+where the pulse frequency and arrival time lead to displacement and an additional complex phase in frequency space, respectively.
+A comparison between the Fourier transform result and the numerical result obtained by solving the differential equations in time domain is shown in Fig.
+S1.
+Figure S1: Comparison between analytical (left, Fourier transform) and numerical (right, obtained by solving the differential equations) results for DT spectrum of a 2LS, plotted in a frame rotating at the central pulse frequency.
+The parameters are $g\sqrt{\mathcal{N}}=3\kappa$, $\tau_{w}=0.05\kappa^{-1}$, $\omega_{p}=\omega_{p^{\prime}}=\omega_{0}$, and $\gamma_{\phi},\gamma=0$.
+S6 Formal integration of pump-probe response
+From Eq.
+(S8), the integrated solution for the vectorized density matrix second order in the pump and first order in the probe, $\vec{\rho}^{(2)(1)}$, is obtained as [denoting the free molecular propagator, i.e., the Green’s function by $\theta(t-t^{\prime})\mathrm{e}^{-\mathrm{i}\mathcal{L}_{0}(t-t^{\prime})}\equiv\mathcal{G}(t-t^{\prime})$]
+$\vec{\rho}^{(2)(1)}(t)=-\mathrm{i}\int_{0}^{t}\mathrm{d}t_{1}\,\mathcal{G}(t-t_{1})\times$ (S21)
+×$\times\left[\mathcal{L}_{\text{int}}^{(2)(1)}(t_{1})\vec{\rho}^{(0)(0)}(t_{1})+\mathcal{L}_{\text{int}}^{(2)(0)}(t_{1})\vec{\rho}^{(0)(1)}(t_{1})\!+\!\mathcal{L}_{\text{int}}^{(1)(1)}(t_{1})\vec{\rho}^{(1)(0)}(t_{1})\!+\!\mathcal{L}_{\text{int}}^{(1)(0)}(t_{1})\vec{\rho}^{(1)(1)}(t_{1})\!+\!\mathcal{L}_{\text{int}}^{(0)(1)}(t_{1})\vec{\rho}^{(2)(0)}(t_{1})\right].$
+Writing out all terms explicitly in terms of the molecular initial state $\vec{\rho}(0)$, yields the following contributions:
+$\vec{\rho}^{(2)(1)}(t)$ $=-\mathrm{i}\int_{0}^{t}\mathrm{d}t_{1}\,\mathcal{G}(t-t_{1})\mathcal{L}_{\text{int}}^{(2)(1)}(t_{1})\vec{\rho}(0)$
+$(-\mathrm{i})^{2}\int_{0}^{t}\mathrm{d}t_{2}\,\int_{0}^{t_{2}}\mathrm{d}t_{1}\,\mathcal{G}(t-t_{2})\mathcal{L}_{\text{int}}^{(1)(1)}(t_{2})\mathcal{G}(t_{2}-t_{1})\mathcal{L}_{\text{int}}^{(1)(0)}(t_{1})\vec{\rho}(0)$
+$(-\mathrm{i})^{2}\int_{0}^{t}\mathrm{d}t_{2}\,\int_{0}^{t_{2}}\mathrm{d}t_{1}\,\mathcal{G}(t-t_{2})\mathcal{L}_{\text{int}}^{(2)(0)}(t_{2})\mathcal{G}(t_{2}-t_{1})\mathcal{L}_{\text{int}}^{(0)(1)}(t_{1})\vec{\rho}(0)$
+$(-\mathrm{i})^{2}\int_{0}^{t}\mathrm{d}t_{2}\,\int_{0}^{t_{2}}\mathrm{d}t_{1}\,\mathcal{G}(t-t_{2})\mathcal{L}_{\text{int}}^{(0)(1)}(t_{2})\mathcal{G}(t_{2}-t_{1})\mathcal{L}_{\text{int}}^{(2)(0)}(t_{1})\vec{\rho}(0)$
+$(-\mathrm{i})^{2}\int_{0}^{t}\mathrm{d}t_{2}\,\int_{0}^{t_{2}}\mathrm{d}t_{1}\,\mathcal{G}(t-t_{2})\mathcal{L}_{\text{int}}^{(1)(0)}(t_{2})\mathcal{G}(t_{2}-t_{1})\mathcal{L}_{\text{int}}^{(1)(1)}(t_{1})\vec{\rho}(0)$
+$(-\mathrm{i})^{3}\int_{0}^{t}\mathrm{d}t_{3}\,\int_{0}^{t_{3}}\mathrm{d}t_{2}\,\int_{0}^{t_{2}}\mathrm{d}t_{1}\,\mathcal{G}(t-t_{3})\mathcal{L}_{\text{int}}^{(1)(0)}(t_{3})\mathcal{G}(t_{3}-t_{2})\mathcal{L}_{\text{int}}^{(1)(0)}(t_{2})\mathcal{G}(t_{2}-t_{1})\mathcal{L}_{\text{int}}^{(0)(1)}(t_{1})\vec{\rho}(0)$
+$(-\mathrm{i})^{3}\int_{0}^{t}\mathrm{d}t_{3}\,\int_{0}^{t_{3}}\mathrm{d}t_{2}\,\int_{0}^{t_{2}}\mathrm{d}t_{1}\,\mathcal{G}(t-t_{3})\mathcal{L}_{\text{int}}^{(1)(0)}(t_{3})\mathcal{G}(t_{3}-t_{2})\mathcal{L}_{\text{int}}^{(0)(1)}(t_{2})\mathcal{G}(t_{2}-t_{1})\mathcal{L}_{\text{int}}^{(1)(0)}(t_{1})\vec{\rho}(0)$
+$(-\mathrm{i})^{3}\int_{0}^{t}\mathrm{d}t_{3}\,\int_{0}^{t_{3}}\mathrm{d}t_{2}\,\int_{0}^{t_{2}}\mathrm{d}t_{1}\,\mathcal{G}(t-t_{3})\mathcal{L}_{\text{int}}^{(0)(1)}(t_{3})\mathcal{G}(t_{3}-t_{2})\mathcal{L}_{\text{int}}^{(1)(0)}(t_{2})\mathcal{G}(t_{2}-t_{1})\mathcal{L}_{\text{int}}^{(1)(0)}(t_{1})\vec{\rho}(0).$ (S22)
+The last term corresponds to the result that would be obtained for conventional free-space spectroscopy with well-separated pulses, i.e., two linear applications of the pump field followed by application of the probe field Mukamel (1995).
+Figure S2: (a) The envelope of the pump and probe pulses in frequency domain is fitted to the lower polariton frequency.
+(b) DT spectrum for excitation of the lower polariton.
+We chose $g\sqrt{\mathcal{N}}=3\kappa$ and a pulse width of $\tau_{w}=\kappa^{-1}$.
+Note that, since we are in a rotating frame, $\omega=0$ corresponds to the central pulse (and therefore the lower polariton) frequency.
+(c) Time evolution of molecular population created by the pump.
+S7 Narrowband spectroscopy of polaritons
+The results presented in the main text were obtained by broadband excitation of the polaritons, i.e., by exciting an (equal) superposition of upper and lower polariton.
+Instead, one may also consider narrowband excitation of the polaritons, by choosing a pulse width in frequency domain $\tau_{w}^{-1}<g\sqrt{\mathcal{N}}$ and fitting the pulse carrier frequency to the polariton frequencies.
+Figs.
+S2(b), (c) show the result obtained for narrowband pump-probe excitation of the lower polariton.
+In this case, no Rabi oscillations are observed since no superposition of upper and lower polariton is excited.
+Narrowband excitation of the upper polariton shows a similar result.
+Figure S3: DT for varying degrees of disorder (a) $\delta=0$, (b) $\delta=0.3\kappa$, and (c) $\delta=0.6\kappa$.
+Left and middle columns show different phase contributions, while the right column presents a cut through the $(2,-1)$ contribution at $\omega=g\sqrt{\mathcal{N}}$ (normalized for comparison across different cases).
+We chose $g\sqrt{\mathcal{N}}=3\kappa$, $\gamma_{\phi}=0.2\kappa$.
+S8 Effects of disorder on DT
+To study the effect of inhomogeneous broadening, we employ a simplified model in which the molecular ensemble is divided into two (equal) subensembles with shifted transition frequencies, $\omega_{\pm}=\omega_{0}\pm\delta$.
+In Fig.
+S3, we illustrate the effect of the degree of disorder $\delta$ on the DT spectra.
+Generally, we find that the DT appears quite robust with increasing disorder.
+At larger degrees of disorder, the dark states are appearing in the spectrum at $\omega=0$, and there is a slight increase in the vacuum Rabi splitting, consistent with previous theoretical treatments Sommer et al.
+(2021); Schwennicke et al.
+(2024); Chen et al.
+(2022).
+Furthermore, the bright state lifetime [as extracted from the DQC $(2,-1)$ contribution] decays faster with increasing disorder.
+S9 Bright and dark states in mean-field theory
+We show here how molecular dark and bright states are arising in the mean-field approach (for a 2LS) Fowler-Wright et al. (2022). The total population is given by
+$p_{\mathrm{tot}}=\sum_{j=1}^{\mathcal{N}}\expectationvalue{\sigma_{j}^{\dagger}\sigma_{j}},$ (S23)
+which in mean field theory becomes
+$p_{\mathrm{tot}}=\mathcal{N}\expectationvalue{\sigma^{\dagger}\sigma}.$ (S24)
+Using that the general population of the $k$-th collective mode is given by $p_{k}=\expectationvalue{\sigma_{k}^{\dagger}\sigma_{k}}$, the population of the bright mode (corresponding to $k=0$) is given by
+$p_{B}=p_{k=0}=\frac{1}{\mathcal{N}}\sum_{j,j^{\prime}=1}^{\mathcal{N}}\expectationvalue{\sigma_{j}^{\dagger}\sigma_{j^{\prime}}^{\phantom{\dagger}}}=\frac{1}{\mathcal{N}}\sum_{j=1}^{\mathcal{N}}\expectationvalue{\sigma_{j}^{\dagger}\sigma_{j}^{\phantom{\dagger}}}+\frac{1}{\mathcal{N}}\sum_{j\neq j^{\prime}}\expectationvalue{\sigma_{j}}^{*}\expectationvalue{\sigma_{j^{\prime}}},$ (S25)
+which in mean field theory becomes
+$p_{B}=\expectationvalue{\sigma^{\dagger}\sigma}+(\mathcal{N}-1)|\expectationvalue{\sigma}|^{2}$ (S26)
+The dark state population $p_{D}=p_{\mathrm{tot}}-p_{B}$ is therefore given
+$p_{D}=(\mathcal{N}-1)\expectationvalue{\sigma^{\dagger}\sigma}-(\mathcal{N}-1)|\expectationvalue{\sigma}|^{2}.$ (S27)
+In the large $\mathcal{N}$ limit, this can be expressed as
+$p_{B}$ $\approx\expectationvalue{\sigma^{\dagger}\sigma}+\mathcal{N}|\expectationvalue{\sigma}|^{2},$ (S28a)
+$p_{D}$ $\approx\mathcal{N}\left(\expectationvalue{\sigma^{\dagger}\sigma}-|\expectationvalue{\sigma}|^{2}\right).$ (S28b)
+Therefore, the bright state corresponds to molecular coherence in the large $\mathcal{N}$ limit.
+S10 Details on phase expansion
+For an arbitrary multilevel system described by a Hamiltonian $\mathcal{H}_{0}$ and a dipole operator $\hat{\mu}=\hat{\mu}^{(+)}+\hat{\mu}^{(-)}$, the mean-field Hamiltonian in the rotating wave approximation is given by
+$\mathcal{H}_{\text{MF}}=\mathcal{H}_{0}+E_{0}\left(\alpha^{*}(t)\hat{\mu}^{(-)}+\alpha(t)\hat{\mu}^{(+)}\right).$ (S29)
+We expand the cavity field and density matrix in terms of the input fields’ amplitudes and phases as
+$\alpha(t)=\sum_{n,m,\vec{v}}\eta_{p}^{n}\eta_{p^{\prime}}^{m}\mathrm{e}^{-\mathrm{i}\vec{v}\cdot\vec{\Phi}}\alpha^{(n)(m)}_{\vec{v}}(t),\quad\rho(t)=\sum_{n,m,\vec{v}}\eta_{p}^{n}\eta_{p^{\prime}}^{m}\mathrm{e}^{-\mathrm{i}\vec{v}\cdot\vec{\Phi}}\rho^{(n)(m)}_{\vec{v}}(t),$ (S30)
+where $\vec{\Phi}=(\Phi_{p},\Phi_{p^{\prime}})^{\top}$ denotes the vector of pump and probe phases and $\vec{v}=(v_{p},v_{p^{\prime}})^{\top}$ denotes the coefficients of the phases, respectively.
+Possible values for the phase coefficients in $(n),(m)$-th order are $v_{p}(n)=\sum_{j=1}^{n}s_{j}$, $v_{p^{\prime}}(m)=\sum_{j=1}^{m}s_{j}$, where each $s_{j}=\pm 1$.
+Note that complex conjugation corresponds to flipping the sign of the phase, i.e., $\left[\alpha^{(n)(m)}_{\vec{v}}\right]^{*}=\alpha^{(n)(m)*}_{-\vec{v}}$.
+The equation for the $(n)(m)$-th order cavity field including phases on the pump and probe input fields reads
+$\dot{\alpha}^{(n)(m)}(t)$ $=-\left(\frac{\kappa}{2}+\mathrm{i}\omega_{c}\right)\alpha^{(n)(m)}-\mathrm{i}E_{0}\mathcal{N}\mathrm{Tr}[\hat{\mu}^{(-)}\rho^{(n)(m)}]-\delta_{1n}\delta_{0m}f_{p}(t-\tau_{p})\mathrm{e}^{-\mathrm{i}\omega_{p}t}\mathrm{e}^{-\mathrm{i}\Phi_{p}}-\delta_{0n}\delta_{1m}f_{p^{\prime}}(t-\tau_{p^{\prime}})\mathrm{e}^{-\mathrm{i}\omega_{p^{\prime}}t}\mathrm{e}^{-\mathrm{i}\Phi_{p^{\prime}}}.$ (S31)
+It is now straightforward to see that, at first order, the only phase contributions of the cavity field which are driven by the input pulses are $(1,0)$ for the field created by the pump and $(0,1)$ for the field created by the probe
+$\dot{\alpha}^{(1)(0)}_{(1,0)}$ $=-\left(\frac{\kappa}{2}+\mathrm{i}\omega_{c}\right)\alpha^{(1)(0)}_{(1,0)}-\mathrm{i}E_{0}\mathcal{N}\mathrm{Tr}[\hat{\mu}^{(-)}\rho^{(1)(0)}_{(1,0)}]-f_{p}(t-\tau_{p})\mathrm{e}^{-\mathrm{i}\omega_{p}t},$ (S32a)
+$\dot{\alpha}^{(0)(1)}_{(0,1)}$ $=-\left(\frac{\kappa}{2}+\mathrm{i}\omega_{c}\right)\alpha^{(0)(1)}_{(0,1)}-\mathrm{i}E_{0}\mathcal{N}\mathrm{Tr}[\hat{\mu}^{(-)}\rho^{(1)(0)}_{(1,0)}]-f_{p^{\prime}}(t-\tau_{p^{\prime}})\mathrm{e}^{-\mathrm{i}\omega_{p^{\prime}}t},$ (S32b)
+while the other possible phase combinations at first order $(-1,0)$ and $(0,-1)$ are contained in the complex conjugate cavity fields.
+All other phase components are not driven at first order and are therefore zero.
+The positive-phase cavity fields can create positive-phase density matrices in first order via the application of $\hat{\mu}^{(+)}$ to a ket (creation of excitation) or to a bra (annihilation of excitation):
+$\dot{\rho}^{(1)(0)}_{(1,0)}$ $=-\mathrm{i}[\mathcal{H}_{0},\rho^{(1)(0)}_{(1,0)}]+\mathcal{D}[\rho^{(1)(0)}_{(1,0)}]-\mathrm{i}E_{0}\alpha^{(1)(0)}_{(1,0)}[\hat{\mu}^{(+)},\rho(0)],$ (S33a)
+$\dot{\rho}^{(0)(1)}_{(0,1)}$ $=-\mathrm{i}[\mathcal{H}_{0},\rho^{(0)(1)}_{(0,1)}]+\mathcal{D}[\rho^{(0)(1)}_{(0,1)}]-\mathrm{i}E_{0}\alpha^{(0)(1)}_{(0,1)}[\hat{\mu}^{(+)},\rho(0)],$ (S33b)
+while the complex conjugate processes correspond to the application of $\hat{\mu}^{(-)}$, thereby imprinting negative phase contributions onto the density matrix.
+Beyond first order, the cavity field is not driven directly anymore.
+Instead, the density matrix is now driven by lower orders of $\alpha$, which determines the possible phase contributions in higher orders.
+The phase combinations for the $(n)(m)$-th order density matrix driven by lower orders of cavity field and density matrix are given by
+$\dot{\rho}^{(n)(m)}_{\vec{v}}=-\mathrm{i}[\mathcal{H}_{0},\rho^{(n)(m)}_{\vec{v}}]+\mathcal{D}[\rho^{(n)(m)}_{\vec{v}}]-\mathrm{i}E_{0}\sum_{j=0}^{n-1}\sum_{j^{\prime}=0}^{m-1}\sum_{\vec{u}+\vec{w}=\vec{v}}\left(\alpha_{\vec{u}}^{(n-j)(m-j^{\prime})}[\hat{\mu}^{(+)},\rho^{(j)(j^{\prime})}_{\vec{w}}]+\alpha_{\vec{u}}^{(n-j)(m-j^{\prime})*}[\hat{\mu}^{(-)},\rho^{(j)(j^{\prime})}_{\vec{w}}]\right),$ (S34)
+where the possible values for $\vec{u}$, $\vec{w}$ are determined by the phase contributions in lower orders. This drives the phase contributions for the cavity field at nonlinear orders $n+m>1$ as
+$\dot{\alpha}^{(n)(m)}_{\vec{v}}(t)=-\left(\frac{\kappa}{2}+\mathrm{i}\omega_{c}\right)\alpha^{(n)(m)}_{\vec{v}}-\mathrm{i}E_{0}\mathcal{N}\mathrm{Tr}[\hat{\mu}^{(-)}\rho^{(n)(m)}_{\vec{v}}].$ (S35)
+S11 Explicit phase expansion for three-level systems
+Let us explicitly consider the phase expansion for a three-level system (3LS), which can, e.g., serve as a simplified model for an anharmonic vibrational mode, described by the dipole operator
+$\hat{\mu}=\mu_{12}\hat{\sigma}_{12}+\mu_{23}\hat{\sigma}_{23}+\mathrm{H.c.},$ (S36)
+where $\hat{\sigma}_{12}=\ket{1}\bra{2}$ | and $\hat{\sigma}_{23}=\ket{2}\bra{3}$ | are the corresponding lowering operators. The mean-field Hamiltonian of a 3LS then reads
+$\mathcal{H}_{\text{MF}}$ $=\omega_{2}\hat{\sigma}_{12}^{\dagger}\hat{\sigma}_{12}^{\phantom{\dagger}}+\omega_{3}\hat{\sigma}_{23}^{\dagger}\hat{\sigma}_{23}^{\phantom{\dagger}}+g_{12}\left(\hat{\sigma}_{12}^{\dagger}\alpha(t)+\mathrm{h.c.}\right)+g_{23}\left(\hat{\sigma}_{23}^{\dagger}\alpha(t)+\mathrm{h.c.}\right),$ (S37)
+with the transition frequencies $\omega_{2}$, $\omega_{3}$ (we set $\omega_{1}=0$), and $g_{12}=\mu_{12}E_{0}$, $g_{23}=\mu_{23}E_{0}$ describe the coupling strengths of the |$\ket{1}$-|$\ket{2}$ and |$\ket{2}$-|$\ket{3}$ transitions to the cavity field, respectively.
+The 2LS limit can simply be obtained by setting $g_{23}=0$.
+In addition, we will consider dephasing affecting the transitions which can be described by the dissipator
+$\mathcal{D}[\rho]=\gamma_{\phi}\left[\hat{\sigma}_{12}^{\dagger}\hat{\sigma}_{12}^{\phantom{\dagger}}\rho\hat{\sigma}_{12}^{\dagger}\hat{\sigma}_{12}-\frac{1}{2}\left\{\rho,\hat{\sigma}_{12}^{\dagger}\hat{\sigma}_{12}^{\phantom{\dagger}}\right\}\right]+\gamma_{\phi}\left[\hat{\sigma}_{23}^{\dagger}\hat{\sigma}_{23}^{\phantom{\dagger}}\rho\hat{\sigma}_{23}^{\dagger}\hat{\sigma}_{23}^{\phantom{\dagger}}-\frac{1}{2}\left\{\rho,\hat{\sigma}_{23}^{\dagger}\hat{\sigma}_{23}^{\phantom{\dagger}}\right\}\right],$ (S38)
+[…]
+First order.
+Assuming identical frequencies for pump and probe $\omega_{p}=\omega_{p^{\prime}}$ and an initial state $\rho(0)=\ket{g}\bra{g}$ |, the equations for the first order coherences in a frame rotating at $\omega_{p}$ are given by
+$\dot{\rho}_{21,(1,0)}^{(1)(0)}$ $=-\left(\frac{\gamma_{\phi}}{2}-\mathrm{i}\Delta_{12}\right){\rho}_{21,(1,0)}^{(1)(0)}-\mathrm{i}g_{12}\alpha^{(1)(0)}_{(1,0)},$ (S39a)
+$\dot{\rho}^{(0)(1)}_{21,(0,1)}$ $=-\left(\frac{\gamma_{\phi}}{2}-\mathrm{i}\Delta_{12}\right)\rho^{(0)(1)}_{21,(0,1)}-\mathrm{i}g_{12}\alpha^{(0)(1)}_{(0,1)},$ (S39b)
+while the equation for the first-order fields are given by
+$\dot{\alpha}_{(1,0)}^{(1)(0)}$ $=-\left(\frac{\kappa}{2}-\mathrm{i}\Delta_{c}\right)\alpha^{(1)(0)}_{(1,0)}-\mathrm{i}\mathcal{N}g_{12}\rho_{21,(1,0)}^{(1)(0)}-f_{p}(t-\tau_{p}),$ (S40a)
+$\dot{\alpha}^{(0)(1)}_{(0,1)}$ $=-\left(\frac{\kappa}{2}-\mathrm{i}\Delta_{c}\right)\alpha_{(0,1)}^{(0)(1)}-\mathrm{i}\mathcal{N}g_{12}\rho_{21,(0,1)}^{(0)(1)}-f_{p^{\prime}}(t-\tau_{p^{\prime}}),$ (S40b)
+where we define the detunings w.r.t.
+the central input pulse frequency $\Delta_{c}=\omega_{p}-\omega_{c}$, $\Delta_{12}=\omega_{p}-\omega_{2}$, and $\Delta_{23}=\omega_{p}-(\omega_{3}-\omega_{2})$.
+The other phase combinations $(-1,0)$ and $(0,-1)$ are contained in the complex conjugate cavity fields.
+Second order.
+The second order populations and coherences are given by
+$\dot{\rho}_{22,(0,0)}^{(2)(0)}$ $=-\mathrm{i}g_{12}\alpha_{(1,0)}^{(1)(0)}\rho_{12,(-1,0)}^{(1)(0)}+\mathrm{i}g_{12}\alpha_{(-1,0)}^{(1)(0)*}\rho_{21,(1,0)}^{(1)(0)},$ (S41a)
+$\dot{\rho}_{22,(1,-1)}^{(1)(1)}$ $=-\mathrm{i}g_{12}\alpha_{(1,0)}^{(1)(0)}\rho_{12,(0,-1)}^{(0)(1)}+\mathrm{i}g_{12}\alpha_{(0,-1)}^{(0)(1)*}\rho_{21,(1,0)}^{(1)(0)},$ (S41b)
+$\dot{\rho}_{31,(2,0)}^{(2)(0)}$ $=-\left[\frac{\gamma_{\phi}}{2}-\mathrm{i}(\Delta_{12}+\Delta_{23})\right]\rho_{31,(2,0)}^{(2)(0)}-\mathrm{i}g_{23}\alpha_{(1,0)}^{(1)(0)}\rho_{21,(1,0)}^{(1)(0)},$ (S41c)
+$\dot{\rho}_{31,(1,1)}^{(1)(1)}$ $=-\left[\frac{\gamma_{\phi}}{2}-\mathrm{i}(\Delta_{12}+\Delta_{23})\right]\rho_{31,(1,1)}^{(1)(1)}-\mathrm{i}g_{23}\alpha_{(0,1)}^{(0)(1)}\rho_{21,(1,0)}^{(1)(0)}-\mathrm{i}g_{23}\alpha^{(1)(0)}_{(1,0)}\rho_{21,(0,1)}^{(0)(1)},$ (S41d)
+while all second-order cavity fields are not driven and therefore remain zero at all times.
+The other phase combinations in second order $(-1,1)$, $(-2,0)$, and $(-1,-1)$ are obtained as the complex conjugates of the above equations.
+Third order.
+The third-order coherences are given by
+$\dot{\rho}_{21,(0,1)}^{(2)(1)}$ $=-\left(\frac{\gamma_{\phi}}{2}-\mathrm{i}\Delta_{12}\right){\rho}_{21,(0,1)}^{(2)(1)}-\mathrm{i}g_{12}\alpha^{(2)(1)}_{(0,1)}+2\mathrm{i}g_{12}\alpha_{(0,1)}^{(0)(1)}\rho_{22,(0,0)}^{(2)(0)}+2\mathrm{i}g_{12}\alpha_{(1,0)}^{(1)(0)}\rho_{22,(-1,1)}^{(1)(1)}-\mathrm{i}g_{23}\alpha^{(1)(0)*}_{(-1,0)}\rho_{31,(1,1)}^{(1)(1)},$ (S42a)
+$\dot{\rho}^{(2)(1)}_{21,(2,-1)}$ $=-\left(\frac{\gamma_{\phi}}{2}-\mathrm{i}\Delta_{12}\right)\rho^{(2)(1)}_{21,(2,-1)}-\mathrm{i}g_{12}\alpha^{(2)(1)}_{(2,-1)}+2\mathrm{i}g_{12}\alpha_{(1,0)}^{(1)(0)}\rho_{22,(1,-1)}^{(1)(1)}-\mathrm{i}g_{23}\alpha^{(0)(1)*}_{(0,-1)}\rho_{31,(2,0)}^{(2)(0)},$ (S42b)
+$\dot{\rho}^{(2)(1)}_{32,(0,1)}$ $=-\left(\gamma_{\phi}-\mathrm{i}\Delta_{23}\right)\rho_{32,(0,1)}^{(2)(1)}-\mathrm{i}g_{23}\alpha^{(0)(1)}_{(0,1)}\rho^{(2)(0)}_{22,(0,0)}-\mathrm{i}g_{23}\alpha^{(1)(0)}_{(1,0)}\rho^{(1)(1)}_{22,(-1,1)}+\mathrm{i}g_{12}\alpha^{(1)(0)*}_{(-1,0)}\rho_{31,(1,1)}^{(1)(1)},$ (S42c)
+$\dot{\rho}^{(2)(1)}_{32,(2,-1)}$ $=-\left(\gamma_{\phi}-\mathrm{i}\Delta_{23}\right)\rho_{32,(2,-1)}^{(2)(1)}-\mathrm{i}g_{23}\alpha^{(1)(0)}_{(1,0)}\rho_{22,(1,-1)}^{(1)(1)}+\mathrm{i}g_{12}\alpha^{(0)(1)*}_{(0,-1)}\rho_{31,(2,0)}^{(2)(0)}.$ (S42d)
+The third-order equations for the fields are given by
+$\dot{\alpha}_{(0,1)}^{(2)(1)}$ $=-\left(\frac{\kappa}{2}-\mathrm{i}\Delta_{c}\right)\alpha_{(0,1)}^{(2)(1)}-\mathrm{i}\mathcal{N}g_{12}\rho_{21,(0,1)}^{(2)(1)}-\mathrm{i}\mathcal{N}g_{23}\rho_{32,(0,1)}^{(2)(1)},$ (S43a)
+$\dot{\alpha}^{(2)(1)}_{(2,-1)}$ $=-\left(\frac{\kappa}{2}-\mathrm{i}\Delta_{c}\right)\alpha_{(2,-1)}^{(2)(1)}-\mathrm{i}\mathcal{N}g_{12}\rho_{21,(2,-1)}^{(2)(1)}-\mathrm{i}\mathcal{N}g_{23}\rho_{32,(2,-1)}^{(2)(1)}.$ (S43b)
+The other phase combinations $(0,-1)$, and $(-2,1)$ are described by the complex conjugate cavity fields.
+The resulting DT spectrum for the different phase contributions is plotted in Fig.
+S4, showing that for a 3LS the $(2,-1)$-contribution also exists for pulse delay times $>(\kappa+\gamma_{\phi})^{-1}$.
+Figure S4: (a) Total DT signal as a function of pulse delay time for a 3LS, as well as phase contributions (b) $(0,1)$, and (c) $(2,-1)$.
+We chose $g_{12}\sqrt{\mathcal{N}}=3\kappa$, $g_{23}=\sqrt{2}g_{12}$, $\omega_{c}=\omega_{p}=\omega_{p^{\prime}}=\omega_{2}$, and $\omega_{3}-\omega_{2}=1.8\omega_{2}$.
+The dephasing was set to $\gamma_{\phi}=0.1\kappa$.
+S12 Extension to multimode cavities
+For many scenarios and applications (e.g., studies of polariton transport, polariton condensation,…), the picture of a single cavity mode does not suffice and instead many photonic modes inside the cavity are required.
+We sketch here how the mean-field treatment can be extended in a straightforward way to also account for such more complicated electromagnetic environments, namely by grouping the molecular system into subsystems Fowler-Wright (2024).
+To this end, we consider a planar cavity which extends a width $W$ perpendicular (transverse) to the cavity axis.
+To enable a tractable numerical treatment, one can coarse-grain the transverse extent into $N_{k}$ discrete points at distance $\Delta r$, $W=N_{k}\Delta r$.
+Assuming the total number of molecules $\mathcal{N}$ inside the cavity to be uniformly distributed along the transverse direction, every one of the grid points is assigned $N_{E}=\mathcal{N}/N_{k}$ molecules interacting with $N_{k}$ photon modes.
+This is described by the Hamiltonian
+$\mathcal{H}=\sum_{k=1}^{N_{k}}\omega_{k}\hat{a}_{k}^{\dagger}\hat{a}_{k}^{\phantom{\dagger}}+\sum_{n=1}^{N_{k}}\sum_{j=1}^{N_{E}}\left[\mathcal{H}_{0}^{nj}+E_{0}\sum_{k=1}^{N_{k}}\left(\hat{a}_{k}\hat{\mu}_{nj}^{(+)}\mathrm{e}^{\mathrm{i}kr_{n}}+\mathrm{h.c.}\right)\right],$ (S44)
+where the index $n=1,\ldots,N_{k}$ labels the position while $j=1,\ldots N_{E}$ labels the particular molecule at a certain position.
+For the cavity, one may assume a quadratic dispersion $\omega_{k}=\omega_{c}+k^{2}c^{2}/(2\omega_{c})$, with $\omega_{c}$ the cavity frequency at $k=0$.
+This parabolic approximation is well‐known to accurately capture the mode structure of planar cavities near normal incidence.
+Within the mean-field approximation, the cavity field is assumed to be classical and all molecules at a position $r_{n}$ are assumed to be identical, i.e., $\mathcal{H}_{0}^{nj}=\mathcal{H}_{0}^{n}$, $\hat{\mu}_{nj}=\hat{\mu}_{n}$, leading to the simplified Hamiltonian
+$\mathcal{H}_{\mathrm{MF}}=\sum_{n=1}^{N_{k}}\mathcal{H}_{0}^{n}+E_{0}\sum_{k=1}^{N_{k}}\sum_{n=1}^{N_{k}}\left(\alpha_{k}(t)\hat{\mu}_{n}^{(+)}\mathrm{e}^{\mathrm{i}kr_{n}}+\mathrm{h.c.}\right).$ (S45)
+The classical evolution of the $N_{k}$ cavity modes driven by pump and probe input pulses with wavevectors $k_{p}$ and $k_{p^{\prime}}$ transverse to the cavity axis is given by
+$\dot{\alpha}_{k}(t)=-\left(\mathrm{i}\omega_{k}+\frac{\kappa}{2}\right)\alpha_{k}(t)-\mathrm{i}E_{0}N_{E}\sum_{n=1}^{N_{k}}\expectationvalue{\hat{\mu}_{n}^{(-)}(t)}\mathrm{e}^{-\mathrm{i}kr_{n}}-\sum_{j=p,p^{\prime}}\eta_{j}\delta_{k_{j},k}f_{j}(t-\tau_{j})\mathrm{e}^{-\mathrm{i}\omega_{j}t},$ (S46)
+where only the modes corresponding to $k=k_{p},k_{p^{\prime}}$ are driven by the input pulses and we assume an equal decay rate $\kappa$ for all cavity modes.
+This set of equations can now, e.g., be used to model pump-probe spectroscopy with pulses entering the cavity at oblique angles.